### v1

# General Riemann Integration on Partially Ordered Net Convergence Structures

Abstract
In this article, we propose a general theory of integration of the Riemann type with respect to arbitrary measures and functions, connected by a continuous bilinear product, with values in abstract Riesz spaces endowed with a general convergence given by nets. This covers both the topological and order based convergences in the literature. We then show that this integral satisfies most of the common order based properties of the objects that comprises integration theory. By establishing a generalized notion of summability on Riesz spaces and an integral build upon countable partitions of the base space and the summability theory, we then shown some uniform convergence theorems for the refereed integrals. An application of these theorems is made to prove various equivalences concerning the Lebesgue, for which we give a brief survey, and Riemann type integrals in partially ordered spaces defined previously. We finish the article with a possible way of classifying general integration procedures defined in abstract convergence structures, and pose some open problems based on them.
keywords:
Bilinear Riemann integral , Net convergence structure , Riesz Spaces , Convergence theorems , Comparison of integrals
[…]
[1]organization=Departamento de Estatística, Universidade Estadual de Campinas, addressline=Rua Sérgio Buarque de Holanda, 651, postcode=13083-970, city=Campinas-SP, country=Brazil
1 Introduction and Historical Context
Since the pioneering work of Riemann [1] on trigonometric functions and integration, a myriad of different types of integrals from the said "Riemann type" have been proposed. Putting aside the much studied Henstock-Kurzweil case, we also have, in the case of partially ordered spaces and with much prominence, the Riemann-Stieltjes integral of [2], [3] and [4], the Birkhoff type integral of [5], the SL, Abstract Derivation Basis and Choquet integrals of [6], [7], [8] and [9], as well as many others that can be gauged from these references and the surveys of [10], [11] and [12]. Therefore, generalized Riemann integration of Henstock-Kurzweil and Choquet types, and the vector-scalar partially ordered case of Birkhoff are well developed from both the general theory, and applications, point of views.
On the other hand, little has been done in generalizing Kolmogorov [13] all encompassing project of integration theory based on Riemann sums to partially ordered spaces. In fact, the author has only been able to localize the works of [14] and [15], which deals with a type of $S^{*}$ integration in convergence groups with examples of the partially ordered case, and the papers of Ionescu-Tulcea [16], [17], [18], which treats the above mentioned Kolmogorov $S^{*}$-integral in the partially ordered Riesz space case, with various examples and simultaneous consideration of the topological case. We also cite the Saks type integrals of [19] and [20] that fit in the general Kolmogorovian framework and will be commented on.
Considering this context and the lack of unification of the integration procedures cited above, we propose, in the present work, to define a general integral based on nets of Riemann sums - using the works of [21], [22], [23], [24] in the topological case - that can be defined in general algebraic structures, equipped with a partial order, and endowed with a specific type of general convergence structure in the sense of [25], called a net convergence structure and recently studied by [26] and [27], but already present in some aspects in [28]. Various additivity and convergence properties of this integration procedure will show that it preserves most of the common properties found in the aforementioned works.
Comparison of this general definition of integral with various types of order convergence based Riemann integrals in the literature (see again [12] for a comprehensive review and bibliography) is also a focus of the present work, and we will show that most of them, as well as Saks type integration procedures, fall withing our framework. Finally, we discuss the equivalence/containment of this general Riemann integral with sequential/net-type approximation Lebesgue type integrals of [29], [30], [31], [32] and others, and propose a general classification and brief bibliographic review of schemes of integration structures available in the literature of partially ordered spaces-valued integration. For information about the topological, algebraic and measure-theoretic concepts that will be used in the context of integration theory in this article, we cite, besides de already given references, [33], [34], [35], [36], [37], [38], [39], [40] and, specially for integration, the more recent publications of [41], [42] and [43].
2 Preliminaries on Notation, Partially Ordered Spaces and Convergence
In in this work, we will, in general, denote partially ordered spaces of any type with uppercase Latin letters $X,Y,Z$, etc. As a rule, all partial orders in the refereed spaces will be denoted by $\leq$, following the custom of classical and recent works on spaces equipped with partial orders, as [35], [36], [44], and [45], as we will not work with more than one partial order, in a fixed space, at a time (therefore, minimizing the possibility of confusion). Differentiation between order notations will be made only when necessary. We will also suppose that every partially ordered space below is Archimedean - for more details, see [36].
As we will be dealing with different types of integrals in this work, we identify them by an superscript consisting of an upper case letter with a direct summary of the name, e.g. $\int^{R}$ for the Riemann integral. The only exception will be the main integral of Section 3, which will be denoted without superscript.
The last point on notation: for a general set $A$, following [46], we denote by $\mathcal{F}(A)$ the collection of all its finite subsets (which is a directed set by inclusion $\subseteq$). We also have the following definition (see, for example, [47]):
Definition 2.1 (Paved Space).
We say that a pair $(\Omega,\mathcal{H})$ consisting of an arbitrary set and a collection of its subsets is a Paved Space if $\emptyset\in\mathcal{H}$.
For the measure and integration results below, we shall need a variety of different convergence notions used in the literature, sometimes not equivalently defined in it, and even not equivalent in general. In this case, comparison of our results with other works should be done considering in detail the definitions below. We usually adopt the most general one to avoid problems of compatibility, and point out equivalences (and differences) wherever possible.
We begin by the following definition given in [30] and [45], which originally is due to Vulikh [35], called $(o_{3})$-convergence:
Definition 2.2 ($(o_{3})$- Convergence of Nets).
Let $X$ be a partially ordered set and $I$, $C$, and $D$ directed sets. We say that a net $\{x_{i}\}_{i\in I}$ in $X$ $(o_{3})$-converges to $x$ in $X$, which we denote by:
$(o_{3})-lim_{i}x_{i}=x,$
if there is an increasing net $\{z_{c}\}_{c\in C}$ and a decreasing net $\{y_{d}\}_{d\in D}$ in $X$ such that:
  1. (i)
    $\sup\{z_{c}:c\in C\}=x=\inf\{y_{d}:d\in D\}$
  2. (ii)
    For every $(c,d)\in C\times D$ there exist $i^{*}\in I$ such that:
    $z_{c}\leq x_{i}\leq y_{d},\ \text{whenever}\ i\geq i^{*}$
Sometimes we may get a simplification by obtaining increasing and decreasing nets in the definition above with same indexing directed set, and then get the following definition of [27] and [26] (see also [45]):
Definition 2.3 (Order Convergence of Nets).
Let $X$ be a partially ordered set and $I$, $C$ directed sets. We say that a net $\{x_{i}\}_{i\in I}$ in $X$ order converges, or $(o)$-converges, to $x$ in $X$, which we denote by:
$(o)-lim_{i}x_{i}=x,$
if there is an increasing net $\{z_{c}\}_{c\in C}$ and a decreasing net $\{y_{c}\}_{d\in C}$ in $X$ such that:
  1. (i)
    $\sup\{z_{c}:c\in C\}=x=\inf\{y_{c}:c\in C\}$
  2. (ii)
    For every $(c,c)\in C\times C$ there exist $i^{*}\in I$ such that:
    $z_{c}\leq x_{i}\leq y_{c},\ \text{whenever}\ i\geq i^{*}$
For the rest of this work, when we say that a net is bounded in a partially ordered space, we mean:
Definition 2.4 (Bounded Net).
Let $X$ be a partially ordered space and $\{x_{i}\}_{i\in I}$ a net in $X$. We say that the net is bounded if and only if there exists an $y\in X$:
$|x_{i}|\leq y,$
for every $i\in I$.
The following characterization of order convergent nets will be useful later (see [48] and [49]):
Proposition 2.1.
Let $X$ be a lattice ordered group. Then, a net $\{x_{i}\}_{i\in I}$ is $(o)$-convergent to $x\in X$ if and only if it is bounded and:
$\limsup_{i}x_{i}=\bigwedge_{i\in I}\bigvee_{i^{{}^{\prime}}\geq i}x_{i^{{}^{\prime}}}=\bigvee_{i\in I}\bigwedge_{i^{{}^{\prime}}\geq i}x_{i^{{}^{\prime}}}=\liminf_{i}x_{i}.$
For lattice ordered groups, we have the following result of [30] and [50], which is also given in [27] and [26] for Riesz spaces:
Proposition 2.2.
Let $X$ be a lattice ordered group and $(I,\alpha)$ a directed set. A net $\{x_{i}\}_{i\in I}$ is order convergent in $X$ if and only if it is $o_{3}$-convergent, which occurs if and only if there is a decreasing net $\{y_{d}\}_{d\in D}$ in $X$ such that $y_{d}\downarrow 0$ and for every $d\in D$ there exists $i^{*}\in I$ such that:
$|x_{i}-x|\leq y_{d}\ ,$
for every $i\geq i^{*}$.
Most of the time we will work with objects less general than lattice ordered groups, e.g. Riesz spaces, in such a way that all of the previous convergence definitions, including the one given in the proposition above, are equivalent. Also, we can show, at least for sequences, that the following result is true (see [45]):
Proposition 2.3.
Let $X$ be a lattice ordered group. A sequence $\{x_{n}\}_{n\in\mathbb{N}}$ in $X$ is $(o_{3}),o$-convergent to $x$ in $X$ if and only if there is a non-increasing sequence $\{u_{n}\}_{n\in\mathbb{N}}$ in $X$ such that $u_{n}\downarrow 0$ and:
$|x_{n}-x|\leq u_{n},$
for every $n\in\mathbb{N}$.
The sequential nature of the result above is essential, as "tail" type convergence results for nets are not usually not true. For more about this subject, see page 184 of [45].
For various proofs in the analytical theory of partially ordered spaces, one hard to work with property of the previous convergences is that $\varepsilon$-type arguments have no analogue. This complicates some proofs related to Cauchy-type conditions and countable sequence and subsequence arguments. To remediate this fact, a different type of convergence, usually spoken as a type of "regulator" or $(D)$-convergence was created, and plays a major role in classical texts, such as [37] and [36], as well as in the vast literature in integration and convergence theory in partially ordered spaces, as given by the surveys of [12], [11], and the recent works of [27] and [26], and their collaborators.
In this case we introduce the following structure associated with $(D)$-convergence, as in [11]:
Definition 2.5 ($(D)$-sequence or Regulator).
Let $X$ be a partially ordered space. We say that a double sequence $\{a_{ij}\}_{(i,j)\in\mathbb{N}\times\mathbb{N}}$ [^1] From now on, except when strictly needed, we will represent a regulator only as $\{a_{ij}\}_{ij}$. in $X$ is a $(D)$-sequence, or regulator, if $a_{ij}\geq a_{ij+1}$ for every $i,n\in\mathbb{N}$ and:
$\bigwedge_{j=1}^{\infty}a_{ij}=0,\ \forall i\in\mathbb{N}$
It follows that every $(D)$-sequence is non-negative: i.e, for every $i,j$, $a_{ij}\geq 0$.
Then, we have the following definition:
Definition 2.6 ((D)-convergence).
Let $X$ be a lattice ordered group and $I$ a directed set. We say that a net $\{x_{i}\}_{i\in I}$ in $X$ is $(D)$-convergent to an element $x\in X$ if there exists a $(D)$-sequence $\{a_{ij}\}$ in $X$ such that for every $\varphi\in\mathbb{N}^{\mathbb{N}}$ there exists $i_{0}\in I$ such that:
$|x_{i}-x|\leq\bigvee_{j=1}^{\infty}a_{j\varphi(j)},$
for every $i\geq i_{0}$.
In general, we have the following relations between this convergence and the previously defined ones (see Chapter 2 of [11]):
Proposition 2.4.
Let $X$ be an lattice ordered group. Then, a $(o)$-convergent sequence $\{x_{n}\}_{n\in\mathbb{N}}$ in $X$ is $(D)$-convergent.
In general, the reverse statement of convergence is not equivalent - see, for example, [36]. The result for nets is also not, in general, true - see again Chapter 2 of [11].
As indicated earlier, $(D)$-convergence can give, in some cases, a substitute for $\varepsilon\setminus 2^{n}$-type arguments in terms of (one version of) the Fremlin Lemma (see [10]):
Theorem 2.1 (Fremlin Lemma).
Let $X$ be a Dedekind $\sigma$-complete weakly $\sigma$-distributive Riesz space. If $\{a_{nij}\}_{nij}$ is a triple sequence of elements of $X$ such that, for each $n\in\mathbb{N}$, $\{a_{nij}\}_{ij}$ is a $(D)$-sequence, then there is a double sequence $\{b_{ij}\}_{ij}$ of elements of $X$ such that, for all $L\in X^{+}\setminus\{0\}$ and $\varphi\in\mathbb{N}^{\mathbb{N}}$,
$L\wedge(\sum_{n=1}^{k}\bigvee_{i=1}^{\infty}a_{ni\varphi(i+n)})\leq\bigvee_{j=1}^{\infty}(L\wedge b_{j\varphi(j)}),$
for all $k\in\mathbb{N}$.
Moreover, $\{S\wedge b_{ij}\}_{ij}$ is a $(D)$-sequence for all $L\in X^{+}\setminus\{0\}$.
A detailed proof of this important result can be found in [51].
We also have the next concept that will play a pivotal role in proving some convergence theorems for the integrals defined in this work. Originally, at least in integration theory, it is due to Matthes [52] and Maitland-Wright [53].
Definition 2.7 (Weak $\sigma$-Distributivity).
Let $X$ be a lattice ordered group. We say that $X$ is weakly $\sigma$-distributive if and only if:
$\bigwedge_{\varphi\in\mathbb{N}^{\mathbb{N}}}(\bigvee_{i=1}^{\infty}a_{i\varphi(i)})=0,$
for every $(D)$-sequence $\{a_{ij}\}$ in $X$.
As an interesting consequence (which we shall use later in the context of integration), we have (see [11]):
Proposition 2.5.
Let $X$ be an Archimedean and weakly $\sigma$-distributive lattice ordered group. Then, a $(D)$-convergent sequence $\{x_{n}\}_{n\in\mathbb{N}}$ in $X$ is $(o)$-convergent.
We finish this introductory section with the next essential definition, which can be found in most books about partially ordered spaces (see, for example, the classic of [35]):
Definition 2.8 (Dedekind Completeness).
Let $X$ be a partially ordered space. We say that $X$ is Dedekind complete if every upper bounded set has a supremum. Equivalently, if every increasing bounded net has an $(o)$-limit (given by its supremum).
For further properties relating to Dedekind completeness, see Chapter 4 of [36].
3 Partially Ordered Net Convergence Structures and Riemann Integration
From the point of view of integration theory, specially considering the breadth and generality of Riemann integration and the various type of convergences in partially ordered spaces defined in the last section, its interesting to consider convergence structures, compatible with the algebraic operations involved, that generalize and unify them - the partially ordered and topological cases being only specific examples of this general structure.
This idea is not new, and actually resurfaced independently in various places in the integration theory literature, such as [14], [12], [40], [54], [15] [55] and [56], all which cites the partially ordered case as the main motivation, or as a particular interesting example. We will leave the development of the general theory of such integration procedures for general spaces, not necessarily partially ordered, for another article, and focus primarily on the case of net convergence structures, in the sense of [27] and [26], with further axioms stated by [40] (see also [12]) and [39], which are tailor made for our work on Riemann integral, in conjunction with the construction of [22] and [24].
We first fix, following the terminology and definitions of [27], the space $\mathcal{N}(X)$ as the set of all admissible nets (see the next comments) in an abstract space $X$ which will always denote, except when specified, a Riesz space - even though most results of this section, and work, are valid for lattice ordered groups or semigroups. It is important to point out that the concept of an admissible net is of a technical nature (based on the axioms of ZFC, as explained in [27]), which we are not going to emphasize much - it is always possible to find, for each net, an admissible net such that the two have equal limit in the sense of the next definition. So, we shall work only with this class (i.e, identify nets with its admissible class) and comment no further.
To give the essential definition, we need (see Page 4 of [27]):
Definition 3.1 (Quasi-Subnet).
Given two nets $\{x_{i}\}_{i\in I}$ and $\{y_{i^{{}^{\prime}}}\}_{I^{{}^{\prime}}}$, we say that the latter is a quasi-subnet of the former if for every $i_{0}\in I$ there exists $i^{{}^{\prime}}_{0}\in I^{{}^{\prime}}$ such that:
$\{y_{i^{{}^{\prime}}}\}_{i^{{}^{\prime}}\geq i^{{}^{\prime}}_{0}}\subseteq\{x_{i}\}_{i\geq i_{0}}$
Then, we have the following main definition from [27]. As in further constructions, $I$ will, in general, denote a general directed set that is to serve as the domain of the nets.
Definition 3.2 (Net Convergence Structure).
A net convergence structure on $X$ is a function
$\eta:X\rightarrow 2^{\mathcal{N}(X)},$
which is defined by the relation $\{x_{i}\}_{i\in I}\in\eta(x)$, written sometimes as the limit notation $\{x_{i}\}_{i\in I}\rightarrow x$, $\lim_{(I,\eta)}x_{i}=x$ or $\lim_{i}x_{i}=x$, such that:
  1. (i)
    If a net $\{x_{i}\}_{i\in I}$ is constant and equals to a $x\in X$, then $\{x_{i}\}_{i\in I}\in\eta(x)$ and its limit is $a$;
  2. (ii)
    If a net converges to $x$, every quasi-subnet of it converges to $x$;
  3. (iii)
    Let $\{z_{i}\}_{i\in I}$ be a net such that $z_{i}\in\{x_{i},y_{i}\}$, for every $i\in I$, with nets $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ such that $\{x_{i}\}_{i\in I}\rightarrow x$ and $\{y_{i}\}_{i\in I}\rightarrow x$. Then $\{z_{i}\}_{i\in I}\rightarrow x$.
For further comparison with other convergences, the following result of [27] is of interest:
Proposition 3.1.
The Axiom 13.2.1 (ii) is equivalent to the following two conditions:
  1. 1.
    If a net converges to $x$, then so does each of its subnets and,
  2. 2.
    If $\{x_{i}\}_{i\in I}$ converges to $x$ and $\{y_{j}\}_{j\in J}$ is tail equivalent to $\{x_{i}\}_{i\in I}$, then $\{y_{j}\}_{j\in J}$ converges to $x$.
Thus, this result shows that this general convergence concept includes the net convergence definition of [28]. It can also be seen that it includes the net convergences of [12], based on axioms of pages 306-307 from it, and [40].
The integral that will be defined below needs only the three axioms of Definition 3.2 of a net convergence structure, but will be supplemented by the following conditions from [12], first appearing in the work of Boccuto and Candeloro [57], which will be used for uniformity with the approaches of the aforementioned authors and integration in partially ordered spaces more generally:
Definition 3.3 (Boccuto-Candeloro Net Convergence Structure).
A Bocutto-Candeloro net convergence structure on (a Riesz space) $X$ is a function
$\eta:X\rightarrow 2^{\mathcal{N}(X)},$
which is defined by the relation $\{x_{i}\}_{i\in I}\in\eta(x)$, written sometimes as the limit notation $\{x_{i}\}_{i\in I}\rightarrow x$, $\lim_{(I,\eta)}x_{i}=x$ or $\lim_{i}x_{i}=x$, such that:
  1. (i)
    If $\{x_{i}\}_{i\in I}\rightarrow x$ and $\{y_{j}\}_{j\in j}\rightarrow y$, then $\{x_{i}+y_{j}\}_{(i,j)\in I\times J}\rightarrow x+y$; if $\{x_{i}\}_{i\in I}\rightarrow x$ and $\{y_{i}\}_{i\in I}\rightarrow y$ are nets such that $x_{i}\leq y_{i}$ for all $i\in I$, or in a cofinal subset of $I$, then $x\leq y$.
  2. (ii)
    If $\{x_{i}\}_{i\in I}\rightarrow x$, then $\{|x_{i}|\}_{i\in I}\rightarrow|x|$.
  3. (iii)
    If $\{x_{i}\}_{i\in I}\rightarrow x$ and $\{y_{i}\}_{i\in I}\rightarrow x$ and $\{z_{i}\}_{i\in I}$ is a net such that there exists $i_{0}\in I$ such that for every $i\geq i_{0}$, $x_{i}\leq z_{i}\leq y_{i}$, then $\{z_{i}\}_{i\in I}\rightarrow x$.
  4. (iv)
    If $\{x_{n}\}_{i\in\mathbb{N}}\rightarrow x$ is a (convergent) sequence, then $\{\frac{1}{n}x_{n}\}_{n\in\mathbb{N}}\rightarrow 0$
  5. (v)
    If a net $\{x_{i}\}_{i\in I}$ is constant and equals to a $x\in X$, then $\{x_{i}\}_{i\in I}\rightarrow a$.
  6. (vi)
    If a net converges to $x$, every quasi-subnet of it converges to $x$.
  7. (vii)
    Let $\{z_{i}\}_{i\in I}$ be a net such that $z_{i}\in\{x_{i},y_{i}\}$, for every $i\in I$, with nets $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ such that $\{x_{i}\}_{i\in I}\rightarrow x$ and $\{y_{i}\}_{i\in I}\rightarrow x$. Then $\{z_{i}\}_{i\in I}\rightarrow x$.
We will also assume, unless stated to the contrary, that all our convergence structures on the Riesz space $X$ will be Hausdorff: that is, each net admits, at most, only one limit. We refer to [27] and [26] for further details and other interesting restrictions.
We will also make use of the following useful definition of a Cauchy type partially ordered convergence (see, for example, [26]):
Definition 3.4 (Cauchy Nets and Complete Net Convergence Structures).
Let $(X,\eta)$ be a Riesz space equipped with a Boccuto-Candeloro net convergent structure $\eta$. We say that a net $\{x_{i}\}_{i\in I}$ is Cauchy if its associated difference
$\{x_{i}-x_{j}\}_{(i,j)\in I\times I},$
is convergent to 0 in the Boccuto-Candeloro net convergence structure $\eta$.
If each Cauchy net in $X$ associated to $\eta$ is convergent [^2] Notice that every convergent sequence is Cauchy in an arbitrary net convergence structure. For details, see [27]., the convergence structure $\eta$ is called Cauchy or, equivalently, $X$ is called complete (in reference to the net convergence structure $\eta$).
Concerning the notions of convergence in partially ordered spaces defined in the previous section, we have, by a consequence of Definition 9.1 and Proposition 9.2 of [27]:
Proposition 3.2 (Equivalence of Convergences in Riesz spaces).
Let $X$ be a partially ordered space. Then, the $(o_{3})$, $(o)$ and $(D)$-convergences are Boccuto-Candeloro net convergence structures. If $X$ is, further, a lattice ordered group, the first two are equivalent.
Further, we have the following result (see [11]):
Proposition 3.3 (Order Convergence is Cauchy).
In a lattice ordered group or Riesz space $X$, order convergence is Cauchy.
Similarly, we have:
Proposition 3.4 ((D)-convergence is Cauchy).
In a lattice ordered group or Riesz space $X$ that is weakly $\sigma$-distributive and Dedekind complete, (D)-convergence is Cauchy.
In other words, a net $\{x_{i}\}_{i\in I}$ in $X$ is $(D)$-convergent if and only if there exists a non-negative bounded double sequence $\{a_{nk}\}_{nk}\subseteq X$ such that $a_{nk}\downarrow 0$ ($k\rightarrow\infty,\forall n\in\mathbb{N}$) and for every $\varphi\in\mathbb{N}^{\mathbb{N}}$ $\exists i_{0}\in I$ such that:
$|x_{i_{1}}-x_{i_{2}}|\leq\bigvee_{n=1}^{\infty}a_{i\varphi(i)},$
for every $i_{1},i_{2}\geq i_{0}$.
As we have not found in the literature a proof for this result for general nets, we supply a proof inspired in the result for integrals in [58].
Proof.
The necessity is clear.
For the sufficiency, let $\{a_{nk}\}_{nk}\subseteq X$ be a non-negative bounded double sequence such that $a_{nk}\downarrow 0$ ($k\rightarrow\infty,\forall n\in\mathbb{N}$) and for every $\varphi\in\mathbb{N}^{\mathbb{N}}$ $\exists i_{0}\in I$ such that:
$|x_{i_{1}}-x_{i_{2}}|\leq\bigvee_{n=1}^{\infty}a_{i\varphi(i)},$
for every $i_{1},i_{2}\geq i_{0}$.
Define the following index set:
$J=\{i\in I:\exists\varphi\in\mathbb{N}^{\mathbb{N}},i=i(\varphi)\}.$
By definition and construction, for $i^{*}\in J$, the set
$\{x_{i}:i\geq i^{*}\},$
is a bounded set, as for $i\geq i^{*}$,
$x_{i}\leq\bigvee_{n=1}^{\infty}a_{nk}+|x_{i^{*}}|,$
which shows the boundedness.
As $X$ is Dedekind complete, there exists the following quantities, following again the notation above:
$a_{i^{*}}=\inf_{i\geq i^{*}}x_{i},$
$b_{i^{*}}=\inf_{i\geq i^{*}}x_{i}.$
Put, then, for arbitrary $i_{1},i_{2}\in J$, $i^{*}=\max(i_{1},i_{2})$ (note that it satisfies the construction above). By construction, $\{x_{i}:i\geq i^{*}\}$ is bounded and:
$a_{i_{1}}=\inf_{i\geq i_{1}}x_{i}\leq\inf_{i\geq i^{*}}x_{i}\leq x_{i}\leq\sup_{i\geq i^{*}}x_{i}\leq\sup_{i\geq i_{2}}x_{i}=b_{i_{2}},$
Therefore,
$\sup_{i\in I}a_{i}\leq\inf_{i\in I}b_{i}.$
Consequently, there exists $x\in X$ such that
$a_{i}\leq y\leq b_{i},$
for every $i\in J$. Now, let $\varphi\in\mathbb{N}^{\mathbb{N}}$ arbitrary. By construction, there exists $i^{*}=i^{*}(\varphi)\in I$ such that
$x_{i_{1}}\leq x_{i_{2}}+\bigvee_{n=1}^{\infty}a_{n\varphi(n)},$
for every $i_{1},i_{2}\geq i^{*}$. Fix such $i_{2}$. Then, taking the supremum with respect to $i_{1}$,
$b_{i^{*}(\varphi)}\leq x_{i_{2}}+\bigvee_{n=1}^{\infty}a_{n\varphi(n)}.$
As a consequence of the weak $\sigma$-distributivity, $\bigwedge_{\varphi\in\mathbb{N}^{\mathbb{N}}}\bigvee_{n=1}^{\infty}a_{n\varphi(n)}=0$, and therefore
$\bigwedge_{\varphi\in\mathbb{N}^{\mathbb{N}}}b_{i^{*}(\varphi)}-\bigvee_{\varphi\in\mathbb{N}^{\mathbb{N}}}a_{i^{*}(\varphi)}=\bigwedge_{\varphi\in\mathbb{N}^{\mathbb{N}}}(b_{i^{*}(\varphi)}-a_{i^{*}(\varphi)})=0.$
On that account,
$x=\bigwedge_{\varphi\in\mathbb{N}^{\mathbb{N}}}b_{i^{*}(\varphi)}=\bigvee_{\varphi\in\mathbb{N}^{\mathbb{N}}}a_{i^{*}(\varphi)}.$
As a consequence, as for every $i\geq i^{*}(\varphi)$,
$x_{i}-x\leq b_{i^{*}(\varphi)}-a_{i^{*}(\varphi)}\leq\bigvee_{n=1}^{\infty}a_{n\varphi(n)}.$
Similarly for the opposing inequality, in such a way that we obtain, finally,
$|x_{i}-x|\leq\bigvee_{n=1}^{\infty}a_{n\varphi(n)},$
which is the convergence of $x_{i}$ to $x$, with respect to $(D)$-convergence, in $X$.
3.1 The Net Riemann Integral
Now, to begin the preparation for integration structures in the context of abstract net convergence of the type above, we need a further notion from the general convergence theory (see [27] and [26]):
Definition 3.5 (Continuous Maps Between Convergent Spaces).
Let $X$ and $Y$ be Riesz spaces with net convergence structures $\eta_{1}$ and $\eta_{2}$ respectively. We say that a map $f:X\rightarrow Y$ is continuous if:
$\{x_{i}\}_{i\in I}\xrightarrow{\eta_{1}}x\implies\{(f(x_{i})\}_{i\in I}\xrightarrow{\eta_{2}}f(x),$
for every net convergent net $\{x_{i}\}_{i\in I}\in\mathcal{N}(X)$.
To integrate measure and functions, we will need, in conjunction to a Boccuto-Candeloro net convergence structures in each image space of these objects, a bilinear structure that connects the two, one on which the measure lives, the other where the functions live, to a third space equipped with a convergence structure of the cited type, with the additional requirement that it preserves order (isotonicity) and continuity of the algebraic operations. More precisely, we suppose the existence of a bilinear "product" as in [30] (see also [38], [11], [59] and [31]):
Definition 3.6 (Product in Riesz spaces).
let $(X,\eta_{1})$, $(Y,\eta_{2})$ and $(Z,\eta_{3})$ be three Riesz spaces with respective Boccuto-Candeloro convergence structures $\eta_{i}$, $i=1,2,3$.
A map
$\bullet:X\times Y\rightarrow Z,$
which we denote by juxtaposition, such that it is:
  1. (i)
    Isotone on the positive cones of $X$ and $Y$, $X^{+}$ and $Y^{+}$ respectively,
  2. (ii)
    Bilinear in each argument,
  3. (iii)
    And such that the partial maps induced by fixing each argument:
    $\bullet_{y}:X\rightarrow Z$
    $\bullet_{x}:Y\rightarrow Z$
    are continuous in the associated Boccuto-Candeloro net convergence structures of the spaces in each case,
is called a bilinear product, or only product.
For the measure-theoretic structures which the integration theory in the present work will be reduced to, and were constructed by [21], [22], [23], [24] in the (Banach or uniform semigroup) topological space-valued case of measures and functions (and products), we use the next construction:
Definition 3.7 (Nets For Integration).
Let $(\Omega,\mathcal{H})$ a fixed paved space. A net:
$\mathfrak{N}=\{(\sigma_{\gamma},\tau_{\gamma})\}_{\gamma\in\Gamma},$
where $\Gamma$ is a directed set and such that:
$\sigma_{\gamma}=\{\sigma_{\gamma_{i}}\}_{i\in I_{\gamma}},$
$\tau_{\gamma}=\{\tau_{\gamma_{i}}\}_{i\in I_{\gamma}},$
are finite families in $\mathcal{H}$ and $\Omega$ respectively, is called a Net for Integration or, in another terminology, a Fleischer-Szabó-Száz net.
Given these two definitions, we have the main structural object of the present section, defined in analogy to Definition 1.5 of Szabó and Száz [22]:
Definition 3.8 (Partially Ordered Integration Structure).
An ordered triple:
$((\Omega,\mathcal{H}),\mathfrak{N},((X,\eta_{1}),(Y,\eta_{2}),(Z,\eta_{3});\bullet)),$
consisting of
  1. (i)
    $(\Omega,\mathcal{H})$, a paved space;
  2. (ii)
    $\mathfrak{N}$, a Fleischer-Szabó-Száz net defined on (i);
  3. (iii)
    $(X,\eta_{1}),(Y,\eta_{2}),(Z,\eta_{3})$, Riesz spaces with Boccuto-Candeloro convergence structures connected by a product $\bullet$;
will be called an partially ordered integration system.
With these definitions, we may state the main object of our integration theory, again borrowing ideas of Szabó and Száz, and Fleischer:
Definition 3.9 (Net Riemann integral).
Let $f:\Omega\rightarrow X$ be a function and $\mu:\mathcal{H}\rightarrow Y$ be a set function, which will be called a measure, both given with respect to a partially ordered integration structure
$((\Omega,\mathcal{H}),\mathfrak{N},((X,\eta_{1}),(Y,\eta_{2}),(Z,\eta_{3});\bullet)).$
Then, defining as:
$S_{\gamma}(f,\mu)=\sum_{i\in I_{\gamma}}f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}}),$
an abstract Riemann sum in $Z$, if the following limit exists,
$\int_{\Omega}fd\mu=\lim_{(\Gamma,\eta_{3})}S_{\gamma}(f,\mu),$
it will be called the Net Riemann integral of $f$ with respect to $\mu$ on $\Omega$, or only the Net Riemann integral of $f$ if the set function is uniquely specified.
For a set $A\in\mathcal{H}$, we say that $f$ is net Riemann integrable on $A$ if:
$\int_{\Omega}f\mathbbm{1}_{A}d\mu,$
exists and in this case define:
$\int_{A}f\mu=\int_{\Omega}f\mathbbm{1}_{A}d\mu$
and note that, in this case, the Riemann sums cannot, in general, be restricted to $\sigma$ such that all its elements are contained in $A$.
We now comment briefly on a possible modification of this integral inspired by [24].
Observation 3.1 (Fleischer [24] Integration Domains and Fleischer Net Riemann Integral).
In general, the integration process above doesn’t take into account specific properties of the set $A\in\mathcal{H}$ in integrating a function $f$ on $A$, in the sense that the Fleischer-Szabó-Száz net of integration is independent of the set $A$. We can refine the integration concept above and ask that, for each $A\in\mathcal{H}$, we have a net in the space of finite collection of sets and points as specified above, defined on a directed set $\Gamma_{A}$. Thus, we obtain a collection of directed sets $\{\Gamma_{A}\}_{A\in\mathcal{H}}$ and Fleischer-Szabó-Száz nets associated with them. The integral of $f$ on $A$ is then the limit along $\Gamma_{A}$ of the Riemann sums of $f$ associated to Fleischer-Szabó-Száz net of $\Gamma_{A}$. This notion of integral will be called the Fleishcer Net Riemann integral in honor of Fleischer [24] contribution and to distinguish it from the construction above.
Questions of the additivity of the set function induced by such integration procedure, as well as subsets integration (i.e, existence on the integral of $f$ on $B$ implies existence on $A$ for each $A\subseteq B$ and $A,B\in\mathcal{H}$), becomes more delicate to answer than for the integral of Definition 3.9. Nevertheless, most basic properties, such as additivity, isotonocity, and others, remain the same for this integral. Therefore, we focus on the basic Net Riemann integral, and offer some comments on Fleischer’s one when appropriate, specially in the last sections.
Following Definition 3.9, we denote by $L^{1}(\mu)$ the space of all functions $f:\Omega\rightarrow X$ Net Riemann integrable on $\Omega$  [^3] Most of the properties below are valid for $f$ integrable on $A\in\mathcal{H}$ with suitable modifications. In general, we also omit the ”on $\Omega$” if it causes no confusion. with respect to the (fixed) set function $\mu:\Omega\rightarrow Y$ in the context of a (fixed) partially ordered integration structure.
Then, with the previous structures in mind, before presenting examples of specific integration procedures that fit in this archetype, we collect general properties of the Net Riemann integral that can be gathered essentially from the definition and without imposing too many conditions on $f$ and/or $\mu$ and its associated spaces:
Theorem 3.1 (Properties of the Net Riemann integral).
The Net Riemann integral has the following properties:
  1. (i)
    Let $f,g\in L^{1}(\mu)$. Then,
    $\int_{\Omega}(f+g)d\mu=\int_{\Omega}fd\mu+\int_{\Omega}gd\mu,$
    i.e, the integral is additive on the space $L^{1}(\mu)$;
  2. (ii)
    Let $f\in L^{1}(\mu)$ such that $f\geq 0$ and $\mu\geq 0$. Then,
    $\int_{\Omega}fd\mu\geq 0,$
    i.e, the integral is a positive map on the space of integrable functions.
  3. (iii)
    Let $\mu\geq 0$ and $f,g\in L^{1}(\mu)$ such that:
    $g\geq f.$
    Then,
    $\int_{\Omega}gd\mu\geq\int_{\Omega}fd\mu,$
    i.e, the integral is an isotone map on $L^{1}(\mu)$ equipped with the partial ordering inherited from $X$.
  4. (iv)
    Let $\mu\geq 0$ and $f\in L^{1}(\mu)$ such that $|f|\in L^{1}(\mu)$, then:
    $|\int_{\Omega}fd\mu|\leq\int_{\Omega}|f|d\mu.$
Proof.
(i) Let
$S_{\gamma}(f,\mu),$
$S_{\gamma}(g,\mu),$
be the Riemann sums of $f$ and $g$ corresponding to the (fixed) integration structure of $L^{1}(\mu)$. Then, it is readily seen that, as the product is bilinear:
$S(f+g,\mu)=\sum_{i\in I_{\gamma}}(f+g)(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})=\sum_{i\in I_{\gamma}}f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})+\sum_{i\in I_{\gamma}}g(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})=S(f,\mu)+S(g,\mu),$
which, by taking the limit with respect to the Boccuto-Candeloro net convergence structure $\eta_{3}$, and using axiom $(i)$ of it, it follows that:
$\lim_{(\Gamma,\eta_{3})}S_{\gamma}((f+g),\mu)=\lim_{(\Gamma,\eta_{3})}S_{\gamma}(f,\mu)+\lim_{(\Gamma,\eta_{3})}S_{\gamma}(g,\mu),$
which is the same as:
$\int_{\Omega}(f+g)d\mu=\int_{\Omega}fd\mu+\int_{\Omega}gd\mu$
(ii) Let $f\in L^{1}$ such that it, and $\mu$, are non-negative (in $X$ and $Y$ respectively). Let the Riemann sum of $f$ in the fixed integration structure given by $S_{\gamma}(f,\mu)$. By the isotone property of the product and the non-negativity of both $f$ and $\mu$, it follows that:
$S_{\gamma}(f,\mu)=\sum_{i\in I_{\gamma}}f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})\geq 0,$
i.e, the sum is non-negative in $Z$ for every $I_{\gamma}$ and $\gamma\in\Gamma$. Therefore, by taking the limit with respect to the Boccuto-Candeloro net convergence structure $\eta_{3}$, and using the second component of axiom $(i)$, and using axiom $(v)$, it follows that:
$\lim_{(\Gamma,\eta_{3})}S_{\gamma}(f,\mu)\geq 0,$
which is equivalent to
$\int_{\Omega}fd\mu\geq 0.$
(iii) Let $f,g\in L^{1}(\mu)$ with corresponding Riemann sums in the fixed convergent structure $S_{\gamma}(f,\mu),S_{\gamma}(g,\mu)$. As the product is isotone, we have that, for every $\gamma\in\Gamma$ and $i\in I_{\gamma}$,
$f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})\geq g(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}}).$
As the bilinear product is isotone, and $Z$ is a Riesz space (therefore, the sum is an isotone map in $Z$ partial ordering), it follows that:
$\sum_{i\in I_{\gamma}}g(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})\geq\sum_{i\in I_{\gamma}}f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}}),$
for every $\gamma\in\Gamma$ and $i\in I_{\gamma}$.
Therefore, by taking the limit with respect to the Boccuto-Candeloro net convergence structure $\eta_{3}$, and using the second part of axiom $(i)$ of it, it follows that:
$\lim_{(\Gamma,\eta_{3})}S_{\gamma}(g,\mu)\geq\lim_{(\Gamma,\eta_{3})}S_{\gamma}(f,\mu),$
which is the same as
$\int_{\Omega}gd\mu\geq\int_{\Omega}fd\mu$
(iv) Follows from the same previous techniques and the use of axioms $(i)$ and $(ii)$ of the Boccuto-Candeloro net convergence structure of $Z$.
We enounce the following direct consequence of Definition 3.9 and Proposition 2.1, which will be useful in more specialized discussions in the next sections:
Corollary 3.1.1.
Suppose $(Z,\eta_{3})$ is a Dedekind Complete Riesz space with $\eta_{3}$ being given by order convergence. Then the integral of a function $f:\Omega\rightarrow X$ with respect to a set function $\mu:\mathcal{H}\rightarrow Y$ exists with respect to $\eta_{3}$ if and only if:
  1. (i)
    The net $\{S_{\gamma}(f,\mu)\}_{\gamma\in\Gamma}$ is bounded in $Z$ and,
  2. (ii)
    The quantities $\limsup_{\gamma}S_{\gamma}(f,\mu)$ and $\liminf_{\gamma}S_{\gamma}(f,\mu)$ both exists, and
    $\limsup_{\gamma}S_{\gamma}(f,\mu)=\bigwedge_{\gamma\in\Gamma}\bigvee_{\gamma^{{}^{\prime}}\geq\gamma}S_{\gamma^{{}^{\prime}}}(f,\mu)=\bigvee_{\gamma\in\Gamma}\bigwedge_{\gamma^{{}^{\prime}}\geq\gamma}S_{\gamma^{{}^{\prime}}}(f,\mu)=\liminf_{\gamma}S_{\gamma}(f,\mu).$
Having proved these basic consequences of the Net Riemann integral, we offer some more clarifying comments on the integral and its properties.
Observation 3.2.
In this generality, some "expected" properties of the Net Riemann integral may not be valid, such as:
  1. (i)
    In general, the Riemann sums $S_{\gamma}(f,\mu)$ will not be monotone, even if $f$ and $\mu$ are positive, in such a way that the integral above cannot be characterized as an infimum/supremum of these nets.
  2. (ii)
    Considering classical, even Lebesgue, integration theory, a common property of the integrals is given by integration of a indicator function:
    $\int_{\Omega}\mathbbm{1}_{A}d\mu=\mu(A),$
    for some notion of integration. In general, for the Net Riemann integral, such relation is not valid. For its validity, more properties have to be asked on $\mu$, which in general are highly restrictive (e.g, supposing that $(\Omega,\mathcal{H})$ is a topological measure space and $\mu$ is outer/inner regular), or are valid only on some specific cases of such general integration procedure. This will be important later.
  3. (iii)
    The continuity properties of such Net Riemann integrals are even more delicate, as the integral is given in a non-Lebesgue (i.e, as a net of integrals of an approximating sequence of (simple, bounded,etc) functions) way, making difficult discussions on the passing of limits inside of the integration sign, something already present is definitions of Riemann integration of Henstock-Kurzweil, $S^{*}$, Sion and others.
  4. (iv)
    The finiteness restriction of the nets defining the integration procedure can be restrictive in some cases, making the definition not amenable to procedures such as the $S^{*}$-integral of [14] which deals with countable sums based on nets of partitions and related points in it.
Having point $(iv)$ of the previous observations in mind, we pass now to a brief discussion of infinite sums in the context of Riesz spaces with Boccuto-Candeloro net convergence structures, the $\sigma$-additivity of positive set functions/measures with values in such spaces, and the remaining definitions of integration procedures used in this work. After these definitions, we provide a possible solution to problem (iv) in relation to the Sion [60] integral in a specific type of convergence.
3.2 Summability in Riesz Spaces and a generalized Kolmogorov Integral
We basically have two notions of converging series, one corresponding to unconditional convergence, and other to conditional convergence, as in the classical reference of Bourbaki [46]. In this context, consider the next definition:
Definition 3.10 (Unconditionally Convergent Series).
Let $(X,\eta)$ be a Riesz space equipped with a Boccuto-Candeloro net convergent structure $\eta$. Then, we say that a net $\{x_{i}\}_{i\in I}$ in $X$, with partial (unordered) sums of elements (in $X$)
$\{\sum_{j\in J}x_{j}\}_{J\in\mathcal{F}(I)},$
has an unconditionally summable (or convergent) series to $x\in X$, which we represent by $x=\sum_{i\in I}x_{i}$, if:
$x=\lim_{(\mathcal{F}(I),\subseteq,\eta)}\sum_{j\in J}x_{j}$
Similarly, we have the notion of conditionally convergence of series, which will be developed only for sequences:
Definition 3.11 (Conditionally Convergent Series).
Let $(X,\eta)$ be a Riesz space equipped with a Boccuto-Candeloro net convergent structure $\eta$. Then, we say that a sequence of elements $\{x_{n}\}_{n\in\mathbb{N}}$, with partial (ordered) sums of elements (in $X$)
$\{\sum_{i=1}^{n}x_{i}\},$
has a conditionally summable (or convergent) series to $x\in X$, which we represent by $x=\sum_{i=1}^{\infty}x_{i}$, if:
$x=\lim_{(\mathbb{N},\eta)}\sum_{i=1}^{n}x_{i}$
Similar notions, specifically the first one, can be found in [28], [61] and [62]. By using a cofinal map (see [28]), or a quasi-subnet argument in the terminology of [27], we may prove the following two results:
Proposition 3.5 (Unconditional Convergence implies Conditional Convergence).
Let $(X,\eta)$ be a Riesz space equipped with a Boccuto-Candeloro net convergent structure $\eta$ and $\{x_{n}\}_{n\in\mathbb{N}}$ a sequence with unconditionally convergent series. Then, it has a conditionally convergent series with same limit.
Regarding to the tail of a series, we have:
Proposition 3.6 (Tail of Series).
Let $(X,\eta)$ be a Riesz space equipped with a Boccuto-Candeloro net convergent structure $\eta$ and $\{x_{n}\}_{n\in\mathbb{N}}$ a sequence with a conditionally convergent series. Then, the tail of the series, i.e., the sequence:
$\{\sum_{i=n}^{\infty}x_{i}\}_{n\in\mathbb{N}},$
is convergent (with respect to $\eta$) to 0.
Both results, specially Proposition 3.5, are interesting in the context of measure theory in partially ordered spaces with convergence structures. Regarding this context, the next definition is also useful:
Definition 3.12 (Monotonically Convergence Property).
Let $(X,\eta)$ be a Riesz space equipped with a Boccuto-Candeloro net convergent structure $\eta$. Then, we say that $(X,\eta)$ has the monotonically convergence property (MCP) if, given an increasing bounded above net with a supremum $\{x_{i}\}_{i\in I}$, it converges to its supremum in $\eta$, i.e.
$\lim_{(I,\eta)}\{x_{i}\}_{i\in I}=\sup_{i\in I}x_{i}$
We have the following classical result, which can be found in various references, such as [30], [50], [36], [12], [11] (for further examples, see specially [12]):
Proposition 3.7 (Spaces with the MCP).
If $X$ is a Dedekind complete Riesz space/lattice ordered group with Boccuto-Candeloro convergence structure $\eta$ being $(o_{3})$ or $(o)$ convergence, then it has the MCP property.
We will not provide a proof, as it is standard and can be found in the above references.
We also have the following example of a topological convergence with the MCP:
Example 3.1.
Let $(X,\eta)$ be the space $L^{0}(\Omega,\Sigma,\mathbb{P})$ of (equivalence classes of) real Borel-measurable functions defined on a probability space, equipped with convergence in probability, and the partial ordering of almost sure pointwise inequality. It is not difficult to check that this convergence is topological and gives a Boccuto-Candeloro net convergence structure in $X=L^{0}(\Omega,\Sigma,\mathbb{P})$. In this case, taking $\{X_{n}\}_{n\in\mathbb{N}}$ a monotone sequence of random variables in $X$ bounded above (by a random variable $Y$), it follows from Theorem 3.5 of [63] that this sequence has a limit in probability given by its supremum. It follows that the present $(X,\eta)$ has the MCP.
In this case, as we will use repeatedly, we have the following result, which can be proved as in Lemma 1.2 of [30]:
Proposition 3.8 (Non-negative Series Convergence).
Let $(X,\eta)$ be a Riesz space equipped with a Boccuto-Candeloro net convergent structure $\eta$ with MCP property and $\{x_{n}\}_{n\in\mathbb{N}}$ a non-negative sequence with conditionally convergent series and partial sums $\{\sum_{i=1}^{n}x_{i}\}_{n\in\mathbb{N}}$ bounded above in $X$. Then, the series associated to $\{x_{n}\}_{n\in\mathbb{N}}$ is unconditionally convergent and:
$\sum_{j=1}^{\infty}x_{i}=\lim_{(\mathcal{F}(\mathbb{N}),\subseteq,\eta)}\sum_{j\in J}x_{j}=\sup_{n\in\mathbb{N}}\sum_{j=1}^{n}x_{i}$
Observation 3.3 (Ideal Infinite Element).
We can dispense with the existence of the supremum, or the MCP property, in the results above by annexing an ideal element $\infty$ to $X$ in such a way that every increasing sequence has a supremum equal to it.
For $(D)$-convergence, which in general has the MCP property only for sequences, we have:
Proposition 3.9.
Let $(X,\eta)$ be a Dedekind complete Riesz space where $\eta$ is the Boccuto-Candeloro net convergence structure given by $(D)$-convergence. Then, a non-negative sequence $\{x_{n}\}_{n\in\mathbb{N}}$ has a conditionally convergent series if and only if it has an unconditionally convergent series with the same limit.
Proof.
The converse is valid is general by Proposition 3.5, so we prove the first implication. As the series of $\{x_{n}\}_{n\in\mathbb{N}}$ is conditionally convergent and $X$ is Dedekind complete, it follows that, by the proposition above and Proposition 2.4:
$\sum_{n=1}^{\infty}x_{n}=(D)-\lim_{n\rightarrow\infty}\sum_{i=1}^{n}x_{i}=(o)-\lim_{n\rightarrow\infty}\sum_{i=1}^{n}x_{i}=\sup_{n\in\mathbb{N}}\sum_{i=1}^{n}x_{i},$
which means that we can find a non-negative sequence $\{p_{n}\}_{n\in\mathbb{N}}$ in $X$ such that $p_{n}\downarrow 0$ and:
$\sup_{n\in\mathbb{N}}\sum_{i=1}^{n}x_{i}-\sum_{i=1}^{n}x_{i}\leq p_{n},$
for each $n$. Define $a_{ij}=p_{j}$, for every $i\in\mathbb{N}$. Then, by the definition of $\{p_{n}\}_{n\in\mathbb{N}}$, $\{a_{ij}\}$ is a $(D)$-sequence. Fix now any $\varphi\in\mathbb{N}^{\mathbb{N}}$ and take, independent of this $\varphi$, $J_{0}=\{1,\cdots,k\}\in\mathcal{F}(\mathbb{N})$, where $k\in\mathbb{N}$, $k>1$, is fixed. Then, if $J\in\mathcal{F}(\mathbb{N})$ is such that $J_{0}\subseteq J$,
$\sum_{j\in J_{0}}x_{j}=\sum_{i=1}^{k}x_{i}\leq\sum_{j\in J}x_{j}.$
Therefore,
$\sup_{n\in\mathbb{N}}\sum_{i=1}^{n}x_{i}-\sum_{j\in J}x_{j}\leq\sup_{n\in\mathbb{N}}\sum_{i=1}^{n}x_{i}-\sum_{j\in J_{0}}x_{j}\leq p_{k}\leq\bigvee_{i=1}^{\infty}a_{i\varphi(i)}=\bigvee_{i=1}^{\infty}p_{\varphi(i)},$
which shows that $\{\sum_{j\in J}x_{j}\}_{J\in\mathcal{F}(\mathbb{N})}$ is $(D)$-convergent to $\sup_{n\in\mathbb{N}}\sum_{i=1}^{n}x_{i}$, and therefore that the sequence has a unconditionally convergent series with the same (conditional) limit.
Usually we deal with series of non-negative measures valued in Riesz spaces, specially in context with integration, for which we now state a fundamental definition:
Definition 3.13 ($\sigma$-additive Measure in Partially Ordered Spaces).
Let $(\Omega,\mathcal{H})$ a paved space and $(Y,\eta)$ be a Riesz space with Boccuto-Candeloro net convergence structure $\eta$. Then, we say that a non-negative set function:
$\mu:\mathcal{H}\rightarrow Y^{+},$
is $\sigma$-additive, or a $\sigma$-additive measure, if for every disjoint sequence $\{A_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{H}$ such that $\bigcup_{n\in\mathbb{N}}A_{n}\in\mathcal{H}$, then
$\mu(\bigcup_{n\in\mathbb{N}}A_{n})=\lim_{(\mathcal{F}(\mathbb{N}),\subseteq,\eta)}\sum_{j\in J}\mu(A_{j})$
Its important to notice that the series in the definition above is a unconditional one, because the use of the (commutative operation of) countable union in its definition - a point forcefully put in the topological context by Flesicher and Traynor [64].
By using Propositions 3.5 and 3.6, its possible to prove the following fact directly:
Proposition 3.10 (Monotone Continuity and $\sigma$-Additivity).
Let $(\Omega,\mathcal{H})$ a paved space by a ring and $(Y,\eta)$ be a Riesz space with Boccuto-Candeloro net convergence structure $\eta$. Then, a non-negative finitely additive set function:
$\mu:\mathcal{H}\rightarrow Y^{+},$
is $\sigma$-additive if and only if for every decreasing sequence (in $\mathcal{H}$) $\{A_{n}\}_{n\in\mathbb{N}}\downarrow\emptyset$,
$\lim_{(\mathbb{N},\eta)}\mu(A_{n})=0,$
where the limit is decreasing in $n$.
It follows that our definition of $\sigma$-additive measure is the same as the one used in the theory of $\sigma$-additive measures in Riesz spaces equipped with order or (D)-convergence - see [12] for details.
Having a notion of series in partially ordered spaces with abstract convergence structures allow us to define a different type of integral that seems different in nature compared to the Net Riemann integral.
Before the definition itself, consider the following auxiliary object:
Definition 3.14 (Choice Function).
Let $\Omega$ be an abstract (i.e, general) set. We say that a function $\delta:2^{\Omega}\rightarrow\Omega$ is a choice function if, for each $A\in 2^{\Omega}$, $\delta(A)=\delta_{A}\in A$.
One more convention: for a countable collection of sets $P=\{\alpha\}_{\alpha\in P}$, we denote for $J\in\mathcal{F}(\mathbb{N})$,
$\sum_{j\in J,\alpha\in P}g(\alpha_{j}),$
where $g$ is a set function, to be the sum of the elements of $P$ indexed by $J$.
Now we may define a new integral, which in its original form is due to Kolmogorov [13], but the modern usage of it is due (initially) to Deleanu [14], and Ionescu-Tulcea in [16] and [17], called the $S^{*}$-integral:
Definition 3.15 (Kolmogorov $S^{*}$-integral).
Given and ordered pair:
$((\Omega,\mathcal{H}),((X,\eta_{1}),(Y,\eta_{2}),(Z,\eta_{3});\bullet)),$
consisting of
  1. (i)
    $(\Omega,\mathcal{H})$, a paved space $\mathcal{H}$;
  2. (ii)
    $(X,\eta_{1}),(Z,\eta_{2})(Z,\eta_{3})$, Riesz spaces with Boccuto-Candeloro convergence structures connected by a product $\bullet$;
we say that a function $f:\Omega\rightarrow X$ is $S^{*}$ integrable on $A\in\mathcal{H}$ with respect to a set function $\mu:\mathcal{H}\rightarrow Y$ if the following iterated limit exists and is independent of each choice function $\delta$:
$\int_{A}^{S^{*}}fd\mu=\lim_{(\mathcal{P}(A),\leq_{ref},\eta_{3})}\lim_{(\mathcal{F}(\mathbb{N}),\subseteq,\eta_{3})}\sum_{j\in J,\alpha\in P}f(\delta_{\alpha_{j}})\mu(\alpha_{j}),$
where $P=\{\alpha_{j}\}_{j\in\mathbb{N}}\in\mathcal{P}(A)$, the set of all disjoint countable coverings of $A$, equipped by the refinement partial ordering $\leq_{ref}$ [^4] Or only $\leq$ if the context is clear., which is assumed non-empty and directed [^5] In general, that requires more hypothesis on the paving $\mathcal{H}$..
In general, we denote the second limit as the series $\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)$.
If in the previous definition we consider $\mathcal{P}(A)$, for $A\in\mathcal{H}$, consisting only of partitions of the set $A$, we call the resulting integral the $S^{*}$-partition Integral, for which the next results and examples continues to be valid.
The integral in Definition 3.15 may be paraphrased as follows:
Definition 3.16 ($S^{*}$-integral Unconditional Series).
With the same structures of Definition 3.15, an map $f$ is $S^{*}$-integrable on $A\in\mathcal{H}$ if and only if the following limit exists:
$\int_{A}^{S^{*}}fd\mu=\lim_{(\mathcal{P}(A),\leq_{ref},\eta_{3})}\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha),$
where the series is understood as in Definition 3.10, and $\mathcal{P}(A)$ is assumed non-empty and directed.
In other words, whenever the $S^{*}$-integral exists, we assume the existence of the series on the integral for partitions at least cofinal for the convergence (i.e, which refines a fixed element of $\mathcal{P}(\Omega)$).
Now point $(iv)$ of Observation 3.2 may be explained: as the Net Riemann integral uses, in its partially ordered integration system, a finite number of sums for each defining net, the $S^{*}$-integral, defined by a limit of an unconditionally convergent series seems not to be contained in it. We will comment about that later on.
For the rest of the text, we denote by $L_{S^{*}}^{1}(\mu)$ the space of all functions $f:\Omega\rightarrow X$ that are $S^{*}$-integrable with respect to the (fixed) set function $\mu:\Omega\rightarrow Y$ on each $A\in\mathcal{H}$ with the integration structures of Definition 3.15.
We now summarize some properties of the $S^{*}$-integral in this abstract setting, most of them given by Deleanu [14]:
Proposition 3.11 (Properties of the $S^{*}$-Integral).
The $S^{*}$-Integral has the following properties:
  1. (i)
    Let $f,g\in L_{S^{*}}^{1}(\mu)$. Then, for each fixed $A\in\mathcal{H}$
    $\int^{S^{*}}_{A}(f+g)d\mu=\int^{S{*}}_{A}fd\mu+\int^{S{*}}_{A}gd\mu,$
    i.e, the integral is additive on the space $L_{S^{*}}^{1}(\mu)$;
  2. (ii)
    Let $f\in L_{S^{*}}^{1}(\mu)$ such that $f\geq 0$ and $\mu\geq 0$. Then, for each fixed $A\in\mathcal{H}$
    $\int^{S^{*}}_{A}fd\mu\geq 0,$
    i.e, the integral is a positive map on the space of integrable functions.
  3. (iii)
    Let $\mu\geq 0$ and $f,g\in L_{S^{*}}^{1}(\mu)$ such that:
    $g\geq f.$
    Then, for each fixed $A\in\mathcal{H}$
    $\int^{S^{*}}_{A}gd\mu\geq\int^{S^{*}}_{A}fd\mu,$
    i.e, the integral is an isotone map on $L_{S^{*}}^{1}(\mu)$ equipped with the partial ordering inherited from $X$.
  4. (iv)
    Let $\mu\geq 0$ and $f\in L_{S^{*}}^{1}(\mu)$ such that $|f|\in L_{S^{*}}^{1}(\mu)$, then for each fixed $A\in\mathcal{H}$:
    $|\int^{S^{*}}_{A}fd\mu|\leq\int^{S^{*}}_{A}|f|d\mu.$
  5. (v)
    Let $f:\Omega\rightarrow X$ and $\mathcal{H}$ an algebra, such that, given $A,B\in\mathcal{H}$ disjoint, $f$ is $S^{*}$-integrable on $A$ and $B$. Then, $f$ is integrable on $A\cup B$ and:
    $\int^{S^{*}}_{A\cup B}fd\mu=\int^{S^{*}}_{A}fd\mu+\int^{S^{*}}_{B}fd\mu$
  6. (vi)
    If $(Z,\eta_{3})$ is complete and $\mathcal{H}$ an algebra, given $A,B\in\mathcal{H}$ such that $A\subseteq B$ in $\mathcal{H}$, $f$ being $S^{*}$-integrable on $B$ implies $S^{*}$-integrability on $A$.
  7. (vii)
    Let $f:\Omega\rightarrow X$ be constant function equal to $c\in X$ and $\mu:\mathcal{H}\rightarrow Y^{+}$ $\sigma$-additive. Then f is $S^{*}$-partition integrable on each $A\in\mathcal{H}$ and:
    $\int_{A}^{S^{*}}fd\mu=c\mu(A)$
For the proof of these results, see [14].
3.3 Examples and Further properties of the (Fleischer) Net Riemann Integral
We now present some examples of the general Net Riemann integral of Definition 3.9. The first three of them can be found, in the topological case and with values in Banach spaces in [21] - its important to notice that all of the examples can be defined in general net convergence monoids, with a partial order or not, but we leave such considerations for a further work.
Notice also that to give examples of the general Net Riemann Integration, and therefore also of the Fleischer Net Riemann integral, is to construct examples of net structures on the (directed) set $\Gamma$ (or $\Gamma_{A}$ for each $A\in\mathcal{H}$ in the case of Fleischer’s integral). For simplicity, we focus on the Net Riemann integral, as the collection of nets for Fleischer Net Convergence integral in Observation 3.1 are obtained directly from the Net Riemann integral case construction. We thus begin with the simplest (abstract) example possible:
Example 3.2 (Kolmogorov S-Integral).
Let $\mathcal{P}^{f}(\Omega)$ be the set of all finite coverings of a set $\Omega$ in the context of a partially ordered integration structure as in Definition 3.9. Suppose also that the paved space in the definition, $(\Omega,\mathcal{H})$, is a ring paved space such that $\Omega$ is a disjoint union of elements of $\mathcal{H}$  [^6] This is to guarantee that $\Gamma$ in this example is non-empty and directed. We can weaken these and get the same results. This comment is also valid for the next examples.. Then, if we define the refinement partial order $\leq$ on $\mathcal{P}^{f}(\Omega)\times\mathcal{F}^{\mathcal{P}}(\Omega)$ as:
$(P,\{t_{i}\}_{i=1}^{m})\leq(Q,\{s_{i}\}_{j=1}^{n})\iff Q\ \text{refines}\ P,$
for $P,Q\in\mathcal{P}^{f}(\Omega)$, $m,n\in\mathbb{N}$, and $t_{i}\in P$ , $s_{j}\in Q$, $(\mathcal{P}^{f}(\Omega)\times\mathcal{F}^{\mathcal{P}}(\Omega),\leq)$ becomes a (downwards) directed set (see [21]).
Then, putting:
$\Gamma=(\mathcal{P}^{f}(\Omega)\times\mathcal{F}(\Omega),\leq),$
and considering the net formed by the identity, we get that the resulting Net Riemann integral is a refinement net limit (with respect to an abstract Bocutto-Candeloro convergence structrure) of finite sums based on the whole covering $P=\sigma\in\Gamma$. This is a generalization (as we consider coverings, not partitions) of the S-integral of Kolmogorov [13].
A non-finite partition version of the following example is given next:
Example 3.3 (Sion Integral).
Considering different types of refinement type integrals, we may also modify Sion [60] integral to our setting. Fix now the notation as in Definition 3.9.
In this setting, consider $\mathcal{D}(\Omega)$ the set formed by all pairs $(P,\Delta)$ such that:
  1. (i)
    $P\in\mathcal{P}(\Omega)$ is, as above, a disjoint covering of $\Omega$.
  2. (ii)
    $\Delta:\mathcal{P}(\Omega)\rightarrow\mathcal{F}(\mathcal{P}(\Omega))$, defined by $\Delta(P)=\mathcal{I}\in\mathcal{F}(P)$, is given, and called a truncation.
Suppose also that the paved space in the definition, $(\Omega,\mathcal{H})$, is a ring-paved space and that $\Omega$ is a disjoint union of elements of $\mathcal{H}$. Now, define a partial order in $\mathcal{D}(\Omega)$, denoted by >>>>, as:
$(Q,\Delta_{2})>>(P,\Delta_{1})\iff Q\ \text{refines}\ P\ \text{and}\ \Delta_{1}(M)\subseteq\Delta_{2}(M),\forall M\in\mathcal{P}(\Omega)\ \text{such that}\ M\ \text{refines}\ Q$
Then, putting:
$\Gamma=\{(P,\tau,\Delta):(P,\Delta)\in\mathcal{D}(\Omega)\ \text{and}\ \tau=\{\tau_{k}\}_{k\in K}\ \text{countable such that}\ \tau_{k}\in\sigma\subseteq P\},$
and considering it equipped with the partial ordering above independent of the chosen points as in Example 3.2 (i.e, the net limit has to be the same for all points or choice functions in the terminology of [21]), it becomes a directed set and the net:
$\mathfrak{N}(P,\tau,\Delta)=(\{\sigma_{i}\}_{i\in\Delta(P)},\{\tau_{i}\}_{i\in\Delta(P)}),$
is a Fleischer-Szabó-Száz net defined on it - for more details, such as the proof that $\Gamma$ is a directed set with the partial orderings above, see [21].
In this case, the Net Riemann integral obtained is the integral from [60] which we call the Sion Integral, which in turn constitutes a generalization of the integrals of Rickart [65] and Phillips [66].
We now give an abstract version of the Henstock-Kurzweil (see [67]) integral as in [68].
Example 3.4 (Száz Gauge Integral).
We now take $\Omega$ to be a topological space and $\mathcal{R}$ the set of all relations $R\subseteq\Omega\times\Omega$ such that $t\in int(R(t))$ for all $t\in\Omega$ and let $\mathcal{H}$ be a family of sets in $\Omega$ such that:
  1. (i)
    For each $A,B\in\mathcal{H}$, there exists a disjoint finite sequence $\{\sigma_{i}\}_{i=1}^{n}$ in $\mathcal{H}$ such that $A\setminus B=\cup_{i=1}^{n}\sigma_{i}$.
  2. (ii)
    For each $A\in\mathcal{H}$ and $R\in\mathcal{R}$, there exists finite sequences $\sigma=\{\sigma_{i}\}_{i=1}^{n}$ and $\tau=\{\tau_{i}\}_{i=1}^{m}$ in $\mathcal{H}$ and $\mathcal{R}$ respectively such that $(\sigma,\tau,R)\in\Gamma$ and $\cup_{i=1}^{n}\sigma_{i}=A$.
Where $\Gamma$ is, in the examples above, the domain of the Fleischer-Szabó-Száz net and is given by the collection of all triples $(\sigma,\tau,R)$ such that $\sigma=\{\sigma_{i}\}_{i\in I}$ and $\tau=\{\tau_{i}\}_{i\in I}$ are finite families in $\mathcal{H}$ and $\Omega$ respectively, and $R\in\mathcal{R}$ is a relation, such that $\sigma$ is a disjoint collection and:
$\tau_{i}\in cl(\sigma_{i})\ \text{and}\ \sigma_{i}\subseteq R(\tau_{i}),$
for all $i\in I$.
We then define a (downwards) directed partial order in $\Gamma$ by: for any $(\sigma,\tau,R)$ and $(\rho,\nu,U)$ in $\Gamma$ we say that
$(\sigma,\tau,R)\leq(\rho,\nu,U),$
if and only if $\sigma$ is a refinement of $\rho$ and $U\subseteq S$.
By the properties $(i)$ and $(ii)$ of the collection $\mathcal{H}$, the directness of $(\Gamma,\leq)$ is a consequence - see [68].
The net defined by these structures:
$\mathfrak{N}(\sigma,\tau,R)=(\sigma,\tau),$
is a Fleischer-Szabó-Száz net which gives the abstract gauge integral of [68].
In general, the relations in $\mathcal{R}$ are given, in specific examples, by gauges (see [67]) in the collection of open sets of the topology of $\Omega$. Thus, the classical gauge integral is contained in this example, as are the integrals of [11], [58], and other integrals of gauge type by specifying the spaces $X,Y,Z$ in Definition 3.9.
We now give an abstract generalization of the construction given in the last example. This is based on the concept of derivation basis in gauge integration, for which our exposition will be based on [69] (see also [7], [6] and the real line case in [70] and [71]). In Henstock [67] and Mcgill [49] terminology, we are going to integrate interval functions on division spaces (therefore, the lattice valued integral of Mcgill [49] is contained in the Net Riemann integral).
Example 3.5 (Derivation Basis Integral).
We consider now the paved space $(\Omega,\mathcal{H})$ equipped with a measure $\mu$ such that $(\Omega,\mathcal{H},\mu)$ is a measure space - i.e, $\mathcal{H}$ is a $\sigma$-algebra and $\mu$ is a real, positive $\sigma$-additive set function.
We call, following [69], a filterbase $\mathcal{B}$ in the product space $\mathcal{I}\times\Omega$, where $\mathcal{I}$ is a family of sets in $\mathcal{H}$ of positive measure, a derivation basis. We assume now the following conditions on $\mathcal{B}$, which we refer to [69] for details:
  1. (i)
    The derivation basis $\mathcal{B}$ ignores no points;
  2. (ii)
    The derivation basis $\mathcal{B}$ has the partitioning property.
In this case, we denote $P_{\beta}(L)$ the set of all $\beta$-partitions of a fixed $\mathcal{B}$-interval $L$. Then, by the partitioning property and the filterbase condition on $\mathcal{B}$, we have that $\{P_{\beta}(L)\}_{\beta\in\mathcal{B}}$ is a filterbase on the product $\mathcal{P}^{f}(L)\times\mathcal{F}(\Omega)$ - [69] again contains all necessary details.
Now, we associate to $\{P_{\beta}(L)\}_{\beta\in\mathcal{B}}$ a Fleischer-Szabó-Száz net. For that, we follow the procedure given in [72] (see also [73] and [74]) for general nets and filterbasis. That is, let
$D(\{P_{\beta}(L)\}_{\beta\in\mathcal{B}})=\{\alpha=(\pi,P_{\beta}(L)):\pi\in P_{\beta}(L)\ \text{and}\ (\pi,P_{\beta}(L))\in\{P_{\beta}(L)\}_{\beta\in\mathcal{B}}\},$
which is directed (as $\{P_{\beta}(L)\}_{\beta\in\mathcal{B}}$ is a filterbase) by the relation:
$(\pi_{1},P_{\beta_{1}}(L))\leq(\pi_{2},P_{\beta_{2}}(L)),$
if and only if $P_{\beta_{2}}(L)\subseteq P_{\beta_{1}}(L)$. Define now the net $\{x_{\gamma}\}_{\gamma\in D(\{P_{\beta}(L)\}_{\beta\in\mathcal{B}})}$ given by:
$\gamma=(\pi,P_{\beta}(L))\mapsto x_{\gamma}=\pi,$
which has, by definition, values in $\mathcal{P}^{f}(L)\times\mathcal{F}(\Omega)$. Then, if we pick, using the same notation as in the examples above,
$\Gamma=D(\{P_{\beta}(L)\}_{\beta\in\mathcal{B}}),$
the net:
$\mathfrak{N}(\pi,P_{\beta}(L))=x_{(\pi,P_{\beta}(L))}=\pi,$
where $\pi=(\sigma_{(\pi,P_{\beta}(L))},\tau_{(\pi,P_{\beta}(L))})$ is, by definition, a finite collection of $\beta$-intervals and points, is a Fleischer-Szabó-Száz net for the set $L$, for which the Riemann sums based on a set function $\eta$ as in Definition 3.9 are given by:
$S_{(\pi,P_{\beta}(L))}(f,\eta)=\sum_{(x,I)\in\pi}f(x)\eta(I),$
which can be seen to be the same as the ones given in Definition 1 of [69] for the integral (in his notation) $\int_{L}\Phi$. Therefore, his integral is contained in the Net Riemann integral.
Finally, we point out that a similar construction taking $\Gamma=\mathcal{B}$ as the (directed set) domain of the Fleischer-Szabó-Száz net can be made. From the point of view of the convergence structures used, it gives the same results as above.
As a last comment for this example, we point out that the by using the inverse construction of associating a net to a filterbase, it should be possible, by taking a specific $\Gamma$, to obtain the equivalence of the two integrals of the present example in this restricted case, but we will not pursue this route here.
We finish this stream of examples with Millington [75] (see also [76]) construction of an abstract integration in terms of runs, an alternative concept to nets and filterbases created by Kenyon and Morse [77] (in fact, with integration theory in mind!). As to not introduce even more concepts in this example, we will translate the results of Millington’s to the language of nets, and also work with a more restricted set of coverings and refer to Millington’s original article for more details concerning the definitions involved here.
Example 3.6 (Millington Process Integral).
For Millington [75] construction, consider $D$ a directed (downwards or upwards) set, $(\Omega,\mathcal{H})$ as in Examples 3.2 and 3.3, $A(\subseteq\Omega)\in\mathcal{H}$ a fixed set, and let $\mathcal{T}$ be the set of all truncations, as in Example 3.3, in $\mathcal{P}(A)$, and $\mathcal{C}$ the set of all choice functions in $\mathcal{H}$ (i.e, $h(C)=c\in C$, for $h:\mathcal{H}\rightarrow\Omega$). Now, consider the net
$\mathfrak{M}:D\times\mathcal{T}\rightarrow\mathcal{P}(A)\times\mathcal{C}\times\mathcal{T},$
defined by:
$\mathfrak{M}_{1}(d,\Delta)=(P_{d},h_{d},\Delta),$
where we endow $D\times\mathcal{T}$ and $\mathcal{P}(A)\times\mathcal{C}\times\mathcal{T}$ with the product directed partial order specific to each case, with the individual directions of the components of each product being the same as in Example 3.3.
Now, with the same structures, consider a second net:
$\mathfrak{M}_{2}:\mathcal{P}(A)\times\mathcal{C}\times\mathcal{T}\rightarrow\mathcal{F}(\mathcal{P}(A))\times\mathcal{F}(\Omega),$
defined by:
$\mathfrak{M}_{2}(P,h,\Delta)=(\Delta(P),h_{\Delta(P)}),$
where $h_{\Delta(P)}$ is the choice function $h$ applied in each member of $\Delta(P)$, i.e, $h_{\Delta(P)}$ is a finite collection of points from $\Omega$.
In this case, take $\Gamma=D\times\mathcal{T}$ with the aforementioned partial directed order. Then, the net (with values in the product of finite collection of sets and points from $\mathcal{H}$ and $\Omega$ respectively):
$\mathfrak{N}:\Gamma\mapsto(\sigma,\tau),$
defined by:
$\mathfrak{N}(d,\Delta)=(\mathfrak{M}_{2}\circ\mathfrak{M}_{1})(d,\Delta),$
that is, $(\sigma,\tau)=(\Delta(P_{d}),h_{d,\Delta(P_{d})})$, is a Fleischer-Szabó-Száz net such that the Riemann sums corresponding to it are equal to the second component of the integration run in Definition 2.3 of [75].
As in the Derivation Basis integral, a suitable choice of a Fleischer-Szabó-Száz net should give rise to a Millington Process integral in general, but we will not pursue this here.
Further examples of the Net Riemann integral, such as the Riemann-Stieltjes integrals of Candeloro [2], Bochner and Ky-Fan [78], and the abstract Riemann integral of Mcshane [54], can be specified and show to be contained in our main integral concept. We leave further details for another work. Some integrals that don’t seem to fit in the general Riemmanian integration schemes presented until now will be pointed out in the last sections of this work.
As a way of obtaining another useful and interesting example, at least in a restricted sense, we now formulate an equivalence result concerning the Net Riemann integral and the $S^{*}$-integral. This aforementioned equivalence is of a restricted nature because it shows that the existence of the $S^{*}$-integral implies the existence of a particular example of the Net Riemann integral when $Z$ equipped with a Boccuto-Candeloro net convergence structure which is topological. In fact, we have:
Proposition 3.12.
Let $f\in L^{1}_{S^{*}}(\mu)$ be a $S^{*}$-integrable function in the context of conditions (i) and (ii) of Definition 3.15, where $Z$ is equipped with a Boccuto-Candeloro net convergence structure which is generated by a vector space topology. Then, $f$ is Sion integrable.
Proof.
As the convergence in $Z$ is topological, we may find a topology $\tau$ for which each convergence net in $Z$ is convergent in this vector topology. Denote by $\mathcal{U}$ the filter of neighborhoods of $\tau$. In this case, as $f$ is $S^{*}$-integrable in $\Omega$ (the proof for other sets is immediate), fix an arbitrary $U\in\mathcal{U}$ and take symmetric and absorbent $V\in\mathcal{U}$ such that $V+V\subseteq U$ and $P_{0}\in\mathcal{P}(\Omega)$ such that, denoting by $s=\int_{\Omega}^{S^{*}}fd\mu$,
$\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)-s\in V,$
for every $P\geq P_{0}$ in $\mathcal{P}(\Omega)$ and independent of the choice function $\delta$.
By definition of the unconditional summability of each series above, for the given $V\in\mathcal{U}$ and $P$ above we may find $J(P)=J_{P}\in\mathcal{F}(\mathbb{N})$ such that, for every $J\in\mathcal{F}(\mathbb{N})$ such that $J_{P}\subseteq J$,
$\sum_{j\in J,\alpha\in P}f(\delta_{\alpha_{j}})\mu(\alpha_{j})-\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)\in V.$
Now, choose $(P_{0},\Delta_{0})\in\mathcal{D}(\Omega)$ such that the truncation is given by:
$\Delta(Q)=\{\{\alpha_{j}\}\}_{j\in J_{Q}},$
where $Q\in\mathcal{P}(\Omega)$ is such that $Q\geq P_{0}$. In this case, for each $(P,\Delta)>>(P_{0},\Delta_{0})$, each sum:
$\sum_{\alpha\in\Delta(P)}f(\delta_{\alpha})\mu(\alpha),$
is of the form $\sum_{j\in J,\alpha\in P}f(\delta_{\alpha_{j}})\mu(\alpha_{j})$, where $J_{P}\subseteq J$. Therefore, by the choice above,
$\sum_{\alpha\in\Delta(P)}f(\delta_{\alpha})\mu(\alpha)\in s+U.$
As the choice of $U\in\mathcal{U}$ is arbitrary, it follows that $f$ is Sion integrable on $\Omega$ and:
$\int_{\Omega}^{S^{*}}fd\mu=s=\int_{\Omega}^{(Si)}fd\mu$
It follows that problem (iv) in Observation 3.2 is solved in the topological case, and we get a new example of the Net Riemann integral with $Z$ equipped with a topological convergence. In particular, we obtain as a "methodological corollary" that the Net Riemann integral can be used effectively in Hewitt-Yosida, Lebesgue and other decompositions for Riesz Spaces - valued set functions, as in [79].
It would be interesting to obtain an answer to the following question regarding the preceding discussion: for which Boccuto-Candeloro structures, besides the topological ones, is Proposition 3.12 true ?
We end this section with some comments on the behavior of the Net Riemann integral on hereditary (i.e, subset) and simple functions integration.
First, we notice that, as indicated before, and commented by Szabó and Száz [22], the usual integration formula for simple functions in point (ii) of Observation 3.2 will not hold for the Net Riemann integral. That can be seen, for example, from the fact that it contains the Henstock-Kurzweil integral in the case of Riesz valued functions and topological scalar measures (see [11]), and even in this case it is necessary to suppose the Radon regularity of the measure to obtain such property. This motivates the next concept, taken from Section 4 of [22]:
Definition 3.17 (Regular Integrators).
We say that a set function $\mu:\mathcal{H}\rightarrow Y$ in the context of the Net Riemann integral in Definition 3.9 is a regular integrator if:
$\int_{\Omega}\mathbbm{1}_{A}d\mu=\mu(A),$
for each $A\in\mathcal{H}$.
It follows directly from (i) of Theorem 3.1 and the properties of the bilinear product in Definition 3.6 that simple functions are then Net Riemann Integrable, that is:
Proposition 3.13 (Integration of Simple Functions).
Let $f:\Omega\rightarrow X$ be a $\mathcal{H}$-simple function, i.e:
$f=\sum_{i=1}^{n}x_{i}\mathbbm{1}_{A_{i}},$
with $\{x_{i}\}_{i=1}^{n}$ a finite sequence of elements of $X$ and $\{A_{i}\}_{i=1}^{n}$ a finite (not necessarily disjoint) sequence of elements of $\mathcal{H}$. Then, if $\mu:\mathcal{H}\rightarrow Y$ is a regular integrator, $f$ is Net Riemann Integrable and:
$\int_{\Omega}fd\mu=\sum_{i=1}^{n}x_{i}\mu(A)$
Lastly, we have the following question of hereditary integration relating to the main integral, i.e: if a function $f$ is Net Riemann integrable on $B\in\mathcal{H}$ and $A\in\mathcal{H}$ is such that $A\subseteq B$, does it implies that $f$ is integrable on $A$ ? As the example of the Riemann-Stieltjes integral of [54] shows, this cannot be positively answered in general.
If $\Gamma$ is, for example, the space of countable coverings/partitions in the case of Kolmogorov’s $S^{*}$-integral, this property can be proved in full generality of net convergence structures, as seen by Theorem 3 of [14]. Similarly for Theorem 53 of [60] on Page 32 and Theorem 2.2 of [75].
At this point it is ideal to comment further the points made in Observation 3.1. For the Fleischer Net Riemann integral, the hereditary integration property, by which we call the existence of integrals for subsets which are contained in integrable ones, seems difficult to obtain without imposing more conditions on the process of associating a Fleischer-Szabó-Száz net to each set $A\in\mathcal{H}$. For specific integrals, such as the ones presented in Examples 3.2-3.6, this result is valid (see the last paragraph).
For our case of general Boccuto-Candeloro net convergence structures and the Fleischer Net Riemann integral, the analogue result of 5°paragraph of [24] seems the best we can get at the moment, which is, in some sense, a result of imposing a Cauchy type condition for the Riemann sums.
Theorem 3.2 (Subset Integration - Fleischer Net Riemann Integral).
Suppose that $Z$ is complete in its net convergence structure. In the context of Fleischer’s modification of Definition 3.9 of the Net Riemann integral in a fixed partially ordered convergence structure, fix $A,B\in\mathcal{H}$ such that $A\subseteq B$. Consider the following hypothesis:
  1. (i)
    For each $\gamma_{B},\gamma_{B}^{{}^{\prime}}\in\Gamma(B)$, there exists $\gamma_{A},\gamma^{{}^{\prime}}_{A}\in\Gamma(A)$ and $\gamma_{A^{C}}\in\Gamma(A^{c})$ such that there exists $\gamma^{1}_{B},\gamma^{2}_{B}\in\Gamma(B)$ such that $\gamma^{1}_{B}\geq\gamma_{B},\gamma^{2}_{B}\geq\gamma^{{}^{\prime}}_{B}$ with corresponding Fleischer-Szabó-Száz nets
    $\gamma^{1}_{B}\rightarrow((\sigma_{\gamma_{A}}\cup\sigma_{\gamma_{A^{c}}}),(\tau_{\gamma_{A}}\cup\tau_{\gamma_{A^{c}}})),$
    $\gamma^{2}_{B}\rightarrow((\sigma_{\gamma^{{}^{\prime}}_{A}}\cup\sigma_{\gamma_{A^{c}}}),(\tau_{\gamma^{{}^{\prime}}_{A}}\cup\tau_{\gamma_{A^{c}}})),$
Then, if $f$ is Fleischer Net Riemann integrable on $B$, it is Net Riemann Integrable on $A$ and:
$\int_{B}fd\mu=\int_{A}fd\mu+\int_{A^{c}}fd\mu$
The strategy of the proof being the same as the already cited results of Theorem 5.3 of [60] on Page 32 and Theorem 2.2 of [75], which are applications of Cauchy nets, we omit it. When a topology or order-type convergence is available, weaker conditions for the validity of the theorem can de found - besides the references above, an archetypic reference would by Proposition 5.2.10 of [10]. Finally, we point out that this result can also be modified to obtain additivity of the set function induced by the Fleischer Net Riemann integral and the Net Riemann integral, as in the analogous case of references [24] and [75], but we will not comment further.
Having finished the results about the fundamental (or "basic") properties of the Net Riemann integral, we now discuss some convergence theorems for the (Fleischer) Net Riemann integral in the context of specific partially ordered net convergence structures.
4 Uniform Convergence Theorems for Net Riemann and $S^{*}$-partition Integrals
To obtain more refined properties of the (Fleischer) Net Riemann integral and correlated concepts introduced in previous sections, we need some convergence theorems, specially (order) uniform ones. The purpose of these convergence theorems are twofold:
  1. (i)
    Obtaining uniform convergence theorems for the Net Riemann integral, and specially the $S^{*}$(-partition) integral allows us to compare them to [30], [32] and to other order based sequential/net-based integrals in the literature (see the next section for a possible list of such integrals) and;
  2. (ii)
    Based on the comparisons contained in item (i), we may obtain monotone convergence theorems and other limit type properties for these integrals, as well as initial steps for a general classification of such integration procedures.
To obtain these convergence theorems, we would like to make as few assumptions as possible on the integrability of the sequence of functions and (order based) limits involved. For that reason, we will adopt a generalization of the uniform convergence theorem of [48] to the Henstock-Kurzweil integral (with respect to the Lebesgue measure) which, compared to [10] - which extends the partially ordered operator-valued framework of [58] - posits less strong integrability conditions on the limit of the approximating functions. Again, we focus mainly on the Net Riemann integral, the modification for the Fleischer case being immediate.
We now fix some conventions for this section:
  1. (i)
    The existence of any integral with values on $Z$ will be understood to take place in the Boccuto-Candeloro net convergence structure given by $(D)$-convergence. All other types of limit will be specified.
  2. (ii)
    The series appearing in the $S^{*}$-partition below will be understood to be taken in order convergence when the issue of existence arises. That will cause no problems with the requirement of $(D)$-convergence in $(i)$, as these will appear only when the functions and measures involved are non-negative and $Z$ weakly $\sigma$-distributive and Dedekind complete: this gives a increasing sequence (or net, in the $\mathcal{F}(\mathbb{N})$ case) of (unordered) sums which, by Propositions 2.5, 3.8 and 3.9 gives a well defined (and equal) limit in $(D)$-convergence as well. Therefore, we have compatibility with Definition 3.15 considering (i), where $\eta_{3}$ is $(D)$-convergence.
We begin by the following useful concept of [48]:
Definition 4.1 (Uniform Convergence in Order).
We say that a sequence of functions $f_{n}:\Omega\rightarrow X$ converges order uniformly to $f:\Omega\rightarrow X$ is there exists a non-negative and non-increasing sequence $\{u_{n}\}_{n\in\mathbb{N}}$ such that $u_{n}\downarrow 0$ and:
$|f_{n}(x)-f(x)|\leq u_{n},$
for every $x\in\Omega$.
One important comment is that we may modify the definition above to make an "tail based" convergence concept (i.e, making the inequality valid for a "sufficient large nn", and not every $n$) based on $p_{n}$, but this modification will not cause any changes in the arguments below, as we can always make "tail convergence" modifications of these (that would not be the case if we consider uniform convergence based on nets of functions).
We also have the following Cauchy definition from [48]:
Definition 4.2 (Cauchy Sequence Uniformly in Order).
We say that a sequence of functions $f_{n}:\Omega\rightarrow X$ is uniformly Cauchy in order, or only uniformly Cauchy, if is there exists a non-negative and non-increasing sequence $\{u_{n}\}_{n\in\mathbb{N}}$ such that $u_{n}\downarrow 0$ and for each $n\in\mathbb{N}$:
$|f_{i}(x)-f_{j}(x)|\leq u_{n},$
for every $x\in\Omega$ and $i,j\geq n$.
It is straightforward to check that every uniformly order convergent sequence of functions if uniformly Cauchy.
For the main convergence theorem, we need the following analogue of Proposition 3.2 of [48]:
Proposition 4.1.
Let $\mu:\mathcal{H}\rightarrow Y^{+}$ be a regular integrator and $\{f_{n}\}_{n\in\mathbb{N}}$ a sequence of Net Riemann Integrable functions, with respect to $\mu$, defined on $\Omega$. If $\{f_{n}\}_{n\in\mathbb{N}}$ is uniformly Cauchy in order, then the following limit exists:
$\limsup_{n\rightarrow\infty}\int_{\Omega}f_{n}d\mu=\bigwedge_{m=1}^{\infty}\bigvee_{i=m}^{\infty}\int_{\Omega}f_{i}d\mu,$
and,
$\int_{\Omega}f_{n}d\mu\xrightarrow[]{o}\bigwedge_{m=1}^{\infty}\bigvee_{i=m}^{\infty}\int_{\Omega}f_{i}d\mu,$
where this convergence is in order.
The proof is the same as in [48], but adapted to our abstract framework. Therefore, to present the modifications and for completeness, we give the full description.
Proof.
By Proposition 2.1, it is sufficient to show that $\{\int_{\Omega}f_{n}d\mu\}_{n\in\mathbb{N}}$ is bounded and that the following holds:
$\bigwedge_{m=1}^{\infty}\bigvee_{i=m}^{\infty}\int_{\Omega}f_{i}d\mu\leq\bigvee_{m=1}^{\infty}\bigwedge_{i=m}^{\infty}\int_{\Omega}f_{i}d\mu,$
as the reverse inequality is always valid for bounded sequences with values in $X$, and the limit in the proposition is a direct consequence of these inequalities.
Now, as $\{f_{n}\}_{n\in\mathbb{N}}$ is uniformly Cauchy, there exists a non-increasing sequence $\{b_{n}\}_{n\in\mathbb{N}}$ consisting of elements of $X$ such that $u_{n}\downarrow 0$ and, for each $n\in\mathbb{N}$,
$|f_{i}(x)-f_{j}(x)|\leq u_{n},\ \forall x\in\Omega,\ i,j\geq k.$
Therefore, given $n\in\mathbb{N}$ and $i,j\geq n$, we obtain from this last equation that:
$f_{i}(x)\leq f_{j}(x)+u_{n},\forall x\in\Omega,$
and, therefore, as these functions are Net Riemann integrable, and the integral in this case is an isotone functional for the given bilinear product, we have that, integrating both sides of this inequality:
$\int_{\Omega}f_{i}d\mu\leq\int_{\Omega}f_{j}d\mu+u_{n}\mu(\Omega).$
In particular, from that it follows that:
$\int_{\Omega}f_{1}d\mu-u_{1}\mu(\Omega)\leq\int_{\Omega}f_{k}d\mu\leq\int_{\Omega}f_{1}d\mu+u_{1}\mu(\Omega),$ (1)
Thus, $\{\int_{\Omega}f_{n}d\mu\}_{n\in\mathbb{N}}$ is a bounded sequence.
By (1), for each $n\in\mathbb{N}$, we have that:
$\bigvee_{i=n}^{\infty}\int_{\Omega}f_{i}d\mu\leq\bigwedge_{j=n}^{\infty}\int_{\Omega}f_{j}d\mu+u_{n}\mu(\Omega),$
which implies that,
$\bigwedge_{m=1}^{\infty}\bigvee_{k=m}^{\infty}\int_{\Omega}f_{k}d\mu-\bigvee_{m=1}^{\infty}\bigwedge_{k=m}^{\infty}\int_{\Omega}f_{k}d\mu\leq\bigvee_{i=n}^{\infty}\int_{\Omega}f_{i}d\mu-\bigwedge_{j=n}^{\infty}\int_{\Omega}f_{j}d\mu\leq u_{n}\mu(\Omega),$
from this last equation, and from the properties of the bilinear product, as $u_{n}\mu(\Omega)\downarrow 0$ when $n\rightarrow\infty$, we get the result.
The main result of this section, which uses the last result, is the analogue of Theorem 3.3 of [48], but modified for the purposes of the Net Riemann integral, which we now state:
Theorem 4.1 (Uniform Converge Theorem for the Net Riemann integral).
Let $Z$ be Dedekind complete and weakly $\sigma$-distributive, $\mu:\mathcal{H}\rightarrow Y^{+}$ a regular integrator and $\{f_{n}\}_{n\in\mathbb{N}}$ a sequence of Net Riemann integrable functions, with respect to $\mu$, defined on $\Omega$. If $f_{n}$ converges order uniformly to $f$, then $f$ is Net Riemann Integrable with respect to $\mu$ and:
$\int_{\Omega}f_{n}d\mu\xrightarrow[]{o}\int_{\Omega}fd\mu,$
where this last convergence is in order in $X$.
Again, in the proof below, we follow mainly [48], with mild modifications for the net Riemann type result. In particular, we will assume that $\Gamma$ is upwards directed, the downwards directed case being a straightforward modification.
Proof.
As $f_{n}$ converges to $f$ order uniformly, there exists a non-increasing sequence $\{u_{n}\}_{k\in\mathbb{N}}$ in $X$ such that $u_{n}\downarrow 0$ and:
$|f_{k}(x)-f(x)|\leq u_{n},\ \forall x\in\Omega\ \text{and}\ k\in\mathbb{N},$ (2)
which implies that $\{f_{n}\}_{n\in\mathbb{N}}$ is Cauchy uniformly.
Therefore, by Proposition 4.1, there exists $L\in Z$ such that the following order limit statement is valid:
$\int_{\Omega}f_{n}d\mu\xrightarrow[]{o}L.$
In other words, there exists a non-decreasing sequence $\{w_{n}\}_{n\in\mathbb{N}}\in Z$ such that $w_{n}\downarrow 0$ and:
$|\int_{\Omega}f_{n}d\mu-L|\leq w_{n},\ \forall n\in\mathbb{N}.$ (3)
As each $f_{n}$ is Net Riemann integrable on $\Omega$, for each $n\in\mathbb{N}$, there exists a $(D)$-sequence $\{a^{(n)}_{ij}\}_{ij}\in Z$ such that, $\forall\ \psi\in\mathbb{N}^{\mathbb{N}}$, exists $\gamma_{n}\in\Gamma$ such that for each $\gamma\geq\gamma_{n}$ in $\Gamma$ we have
$|\sum_{i\in I_{\gamma}}f_{n}(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})-\int_{\Omega}f_{n}d\mu|\leq\bigvee_{i=1}^{\infty}a^{(n)}_{i\psi(i)}$ (4)
Therefore for $\gamma\geq\gamma_{n}$, using (2),(3) and (4), the following inequalities are valid:
$|\sum_{i\in I_{\gamma}}f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})-L|\leq|\sum_{i\in I_{\gamma}}f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})-\sum_{i\in I_{\gamma}}f_{n}(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})|$
$|\sum_{i\in I_{\gamma}}f_{n}(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})-\int_{\Omega}f_{n}d\mu|+|\int_{\Omega}f_{n}d\mu-L|$
$\leq|\sum_{i\in I_{\gamma}}(f(\tau_{\gamma_{i}})-f_{n}(\tau_{\gamma_{i}}))\mu(\sigma_{\gamma_{i}})|+\bigvee_{i=1}^{\infty}a_{i\phi(i)}+w_{n}.$
Continuing from this last inequality we get that it is
$\leq\sum_{\alpha\in P}|f(\tau_{\gamma_{i}})-f_{n}(\tau_{\gamma_{i}})|\mu(\sigma_{\gamma_{i}})+\bigvee_{i=1}^{\infty}a^{(n)}_{i\psi(i)}+w_{n}$
$\leq\sum_{i\in I_{\gamma}}u_{n}\mu(\sigma_{\gamma_{i}})+\bigvee_{i=1}^{\infty}a^{(n)}_{i\psi(i)}+w_{n}=u_{n}\mu(\Omega)+\bigvee_{i=1}^{\infty}a^{(n)}_{i\psi(i)}+w_{n}$
In resume, until now we have the following inequality:
$|\sum_{i\in I_{\gamma}}f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})-L|\leq u_{n}\mu(\Omega)+w_{n}+\bigvee_{i=1}^{\infty}a^{(n)}_{i\psi(i)},\ \forall\ \gamma\geq\gamma_{n}.$ (5)
Now, consider the following equation defining the bounded triple sequence $\{c_{nij}\}_{nij}$ in $Z$:
$c_{nij}=\begin{cases}w_{j}+u_{j}\mu(S),&\text{if $n=1$},\\ a^{(n-1)}_{ij},&\text{if $n\geq 2,$}\end{cases}$
for $i,j\in\mathbb{N}$.
In this case, notice that, for each $n\in\mathbb{N}$ fixed, the bounded double sequence $\{c_{nij}\}_{ij}$ is a $(D)$-sequence of elements of $Z$ because, for each $i\in\mathbb{N}$, $\{c_{nij}\}_{j\in\mathbb{N}}$ is non-increasing and $c_{nij}\downarrow 0$, when $j\rightarrow\infty$.
Now, let $L^{{}^{\prime}}=u_{1}\mu(\Omega)+w_{1}+\bigvee_{i,j=1}^{\infty}a^{(1)}_{ij}$. If $L^{{}^{\prime}}=0$, then $u_{1}=0$ and, by (2), $f=f_{k}$ $\forall k\in\mathbb{N}$. Assuming, then, that $L^{{}^{\prime}}\in Z^{+}\setminus\{0\}$ exists, by Fremlin Lemma, there exists a double sequence $\{b_{ij}\}_{ij}$ in $Z$ such that $\{L^{{}^{\prime}}\wedge b_{ij}\}_{ij}$ is a $(D)$-sequence and, $\forall\psi\in\mathbb{N}^{\mathbb{N}}$,
$L^{{}^{\prime}}\wedge(\sum_{r=1}^{n}\bigvee_{i=1}^{\infty}c_{ri\psi(i+r)})\leq\bigvee_{j=1}^{\infty}(L^{{}^{\prime}}\wedge b_{j}\psi(j)),\ \forall n\in\mathbb{N}.$ (6)
We now show, that, the $(D)$-sequence $\{a_{ij}\}_{ij}$ given by:
$a_{ij}=(L^{{}^{\prime}}\wedge b_{ij}),$
is the $(D)$-sequence for the existence of the Net Riemann integral of $f$.
In fact, given $\varphi\in\mathbb{N}^{\mathbb{N}}$, let $p=\min_{i\in\mathbb{N}}\varphi(i+1)$ and $\psi\in\mathbb{N}^{\mathbb{N}}$ the function defined by $\psi(j)=\varphi(j+p+1)$, $\forall j\in\mathbb{N}$. By (5), for $\gamma\geq\gamma_{p}$, $\gamma_{p}=\gamma_{p}(\phi)$, all in $\Gamma$, we have that:
$|\sum_{i\in I_{\gamma}}f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})-L|\leq u_{p}\mu(\Omega)+w_{p}+\bigvee_{i=1}^{\infty}a^{(p)}_{i\varphi(i+p+1)}.$ (7)
Also, we have:
$\bigvee_{j=1}^{\infty}(u_{\varphi(j+1)}\mu(\Omega)+w_{\varphi(j+1)})=u_{p}\mu(\Omega)+w_{p}.$
Therefore, from (7), we may write that:
$|\sum_{i\in I_{\gamma}}f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})-L|\leq\bigvee_{j=1}^{\infty}(u_{\varphi(j+1)}\mu(\Omega)+w_{\varphi(j+1)})+\bigvee_{i=1}^{\infty}a^{(p)}_{i\varphi(i+p+1)},$
that is,
$|\sum_{i\in I_{\gamma}}f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})-L|\leq\bigvee_{j=1}^{\infty}c_{ij\varphi(j+1)}+\bigvee_{i=1}^{\infty}c_{(p+1)i\varphi(i+p+1)}\leq\sum_{r=1}^{p+1}\bigvee_{i=1}^{\infty}c_{ri\phi(i+r)}.$ (8)
On the other hand, by (5), and for $n=1$, we have that:
$|\sum_{i\in I_{\gamma}}f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})-L|\leq u_{1}\mu(S)+w_{1}+\bigvee_{i,j=1}^{\infty}a^{(1)}_{ij}=L^{{}^{\prime}},$ (9)
whenever $\gamma\geq\gamma_{1}$, with, as above, $\gamma_{1}\in\Gamma$.
Now, taking $\gamma_{0}$ an upper bound of $\gamma_{1}$ and $\gamma_{p}$ in $\Gamma$ (which always exists, as $\Gamma$ is upwards directed), for $\gamma\geq\gamma_{0}$, we have that, by (6), (8) and (9),
$|\sum_{i\in I_{\gamma}}f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})-L|\leq L^{{}^{\prime}}\wedge(\sum_{r=1}^{p+1}\bigvee_{i=1}^{\infty}c_{ri\varphi(i+r)})\leq\bigvee_{j=1}^{\infty}(L^{{}^{\prime}}\wedge b_{j\varphi(j)})=\bigvee_{j=1}^{\infty}a_{j\varphi(j)},$
which shows that, as indicated before, that $\{a_{ij}\}_{ij}$ is indeed the $(D)$-sequence for the existence of the Net Riemann integral of $f$, implying that $f$ is Net Riemann integrable and, by an appeal to Proposition 2.5:
$\int_{\Omega}f_{n}d\mu\xrightarrow[]{o}\int_{\Omega}fd\mu,$
This finishes the proof of the theorem.
4.1 Null sets and a Uniform Convergence Theorem for the S*-partition Integral
For this subsection, we will now present an uniform convergence for the $S^{*}$-partition integral outside of (to be defined) nulls sets of $\mu$. This will be used in an application bellow about the comparison of Riemann with Lebesgue type net based integrals In general, this type of result will not be true for the general Net Riemann integral which contains, in its special cases, integrals which behaves quite badly with respect to null sets integration - see, for exemple, the Henstock-Kurzweil case in [11]. Nevertheless, we also present such a theorem for the specific case of the Sion integral.
We begin wit the following classical definition.
Definition 4.3 (Null Set and A.E. Properties).
Given a set function $\mu:\mathcal{H}\rightarrow Y$, we say that a set $A\in\mathcal{H}$ is $\mu$-null if $\mu(N)=0$.
We then say that some property (based on $\Omega$) is valid $\mu-a.e.$ if there is some null set $N\in\mathcal{H}$ such that the property is valid outside $N$, i.e in $\Omega\setminus N$.
Before presenting the next convergence results, we prove three minor results.
Proposition 4.2.
Let $\mu:\mathcal{H}\rightarrow Y^{+}$ be a positive additive set function. If $B\in\mathcal{H}$ is $\mu$-null and $\subseteq B, AA\in\mathcal{H}$, then $B$ is $\mu$-null
Proof.
In fact, we have that, by the non-negativity of $\mu$,
$\mu(A)\leq\mu(B),$
which shows, as $\mu(B)=0$ and $\mu(A)\geq 0$, that $\mu(A)=0$. ∎
Also, we shall need the next result for integration on null sets.
Proposition 4.3.
Let $\mu:\mathcal{H}\rightarrow Y^{+}$, $N$ a $\mu-null$ set and $f:\Omega\rightarrow X$ a function. Then, $f$ is $S^{*}$-partition integrable on $N$ and:
$\int_{N}^{S^{*}}fd\mu=0$
Proof.
Let $P\in\mathcal{P}(N)$ be any partition of $N$ in $\mathcal{H}$-measurable sets and $\delta$ an arbitrary choice function. By the proposition above, we have the relation
$\mu(\alpha)=0,\ \forall\alpha\in P,$
and, as a consequence, we have:
$\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)=0.$
Therefore, considering an arbitrary $(D)$-sequence $\{a^{(n)}_{ij}\}_{ij}\in Z$ we have that, $\forall\ \psi\in\mathbb{N}^{\mathbb{N}}$, there exists $P\in\mathcal{P}(N)$ such that for each $P\geq P_{n}$ in $\mathcal{P}(N)$ the unconditionally convergent series of $f$ and $\mu$ exists in $P$ and:
$|\sum_{\alpha\in P}f_{n}(\delta_{\alpha})\mu(\alpha)-0|\leq\bigvee_{i=1}^{\infty}a^{(n)}_{i\psi(i)},$
which shows that $f$ is $S^{*}$-partition integrable on $N$ and its integral is 0. ∎
As a corollary that we will use frequently below (without mention) is:
Corollary 4.1.1.
Let $\mu:\mathcal{H}\rightarrow Y^{+}$, $N$ a $\mu-null$ set and $f:\Omega\rightarrow X$ a $S^{*}$-partition integrable function on $\Omega$. Then,
$\int_{\Omega}^{S^{*}}fd\mu=\int_{\Omega\setminus N}^{S^{*}}f\mu.$
Proof.
As $f$ is integrable on $\Omega$, by (vi) of Theorem 3.11, we have that it is integrable on $\Omega\setminus N$ and, as $f$ is integrable on $N$ by the last result:
$\int^{S^{*}}_{\Omega}fd\mu=\int^{S^{*}}_{N}fd\mu+\int^{S^{*}}_{\Omega\setminus N}fd\mu.$
As $\int_{N}^{S{*}}fd\mu=0$ by the last cited result, if follows that:
$\int^{S^{*}}_{\Omega}fd\mu=\int^{S^{*}}_{\Omega\setminus N}fd\mu$
Later, we will prove an stronger result concerning $S^{*}$-partition integration outside of a null set.
We now formulate the following modification of uniform convergence in order:
Definition 4.4 (Almost Everywhere Uniform Convergence).
We say that a sequence of functions $f_{n}:\Omega\rightarrow X$ converges uniformly almost everywhere in order on $\Omega$ to a function $f:\Omega\rightarrow X$, written as $(u)-\lim_{n\rightarrow\infty}f_{n}=f\ \mu-\text{a.e.}$, if there exists a null set $N\in\mathcal{H}$ such that $f_{n}$ converges order uniformly to $f$ on $\Omega\setminus N$.
In this case, we have the following modifications of Proposition 4.1 and Theorem 4.1:
Proposition 4.4.
Let $X$ be a Dedekind $\sigma$-complete Riesz space that is weakly $\sigma$-distributive and $\{f_{n}\}_{n\in\mathbb{N}}$ a sequence of $S^{*}$-partition integrable functions on $\Omega$. Suppose that $\mu:\mathcal{H}\rightarrow Y^{+}$ is $\sigma$-additive. If $\{f_{n}\}_{n\in\mathbb{N}}$ is Cauchy uniformly almost everywhere in order, then the following limit exists:
$\limsup_{n\rightarrow\infty}\int^{S^{*}}_{\Omega\setminus N}f_{n}d\mu=\bigwedge_{m=1}^{\infty}\bigvee_{i=m}^{\infty}\int^{S^{*}}_{\Omega\setminus N}f_{i}d\mu,$
and,
$\int^{S^{*}}_{\Omega}f_{n}d\mu\xrightarrow[]{o}\bigwedge_{m=1}^{\infty}\bigvee_{i=m}^{\infty}\int^{S^{*}}_{\Omega\setminus N}f_{i}d\mu,$
where $N$ is the null set in the uniformly almost everywhere convergence in order related to $f_{n},f$.
Theorem 4.2 (Almost Everywhere Uniform Converge Theorem for the $S^{*}$-partition Integral).
Let $X$ be a Dedekind complete Riesz space that is weakly $\sigma$-distributive and $\{f_{n}\}_{n\in\mathbb{N}}$ a sequence of non-negative $S^{*}$-partition integrable functions in $S$. Suppose that $\mu:\mathcal{H}\rightarrow Y^{+}$ is $\sigma$-additive. If $f_{n}$ converges uniformly almost everywhere in order to $f$, a non-negative function, then $f$ is $S^{*}$-partition integrable on $\Omega\setminus N$ and:
$\int^{S^{*}}_{\Omega}f_{n}d\mu\xrightarrow[]{o}\int^{S^{*}}_{\Omega\setminus N}fd\mu,$
where $N$ is the null set in the uniformly almost everywhere convergence in order related to $f_{n},f$.
The proof of Proposition 4.4 is the same for the original one, and requires only the modification of including the null set, which doesn’t affect the proof. Therefore, we omit it.
For Theorem 4.2, we need some small, but important modifications of Theorem 4.1 relating to the nature of the $S^{*}$-partition integral. With the danger of being a little bit repetitive, we present the proof and the necessary modifications.
Proof.
First, we prove the infinite summations that occurs in the definition of the $S^{*}$-partition integral of $f$, on $\Omega\setminus N$ exist. Throughout the proof, consider $\delta$ an arbitrary choice function.
As $f_{n}$ converges to $f$ uniformly almost everywhere in order, there exists a non-increasing sequence $\{u_{n}\}_{n\in\mathbb{N}}$ such that $u_{n}\downarrow 0$ and:
$|f_{n}(x)-f(x)|\leq u_{n},\ \forall x\in\Omega\setminus N\ \text{and}\ n\in\mathbb{N}.$
Fix an arbitrary $n\in\mathbb{N}$ such that the inequality above is valid, and let $P\in\mathcal{P}(\Omega\setminus N)$ a partition of in the definition of the $S^{*}$–partition integral of $f_{n}$ (as $f_{n}$ is integrable on $\Omega$, it is on $\Omega\setminus N$ by the results above) such that the following (increasing limit) exists:
$\sum_{\alpha\in P}f_{n}(\delta_{\alpha})\mu(\alpha),$
and consider an arbitrary, but fixed, $J\in\mathcal{F}(\mathbb{N})$. Then, we have that, for these choices:
$|\sum_{j\in J,\alpha\in P}f(\delta_{\alpha_{j}})\mu(\alpha_{j})|\leq\sum_{j\in J,\alpha\in P}|f(\delta_{\alpha_{j}})-f_{n}(\delta_{\alpha_{j}})|\mu(\alpha_{j})+\sum_{j\in J,\alpha\in P}f_{n}(\delta_{\alpha_{j}})\mu(\alpha_{j}),$
but, from the first inequality above,
$\sum_{j\in J,\alpha\in P}|f(\delta_{\alpha_{j}})-f_{n}(\delta_{\alpha_{j}})|\mu(\alpha_{j})\leq u_{n}\mu(\Omega\setminus N),$
and,
$\sum_{j\in J,\alpha\in P}f_{n}(\delta_{\alpha_{j}})\mu(\alpha_{j})\leq\sum_{\alpha\in P}f_{n}(\delta_{\alpha})\mu(\alpha),$
which implies that, by the previous inequalities:
$|\sum_{j\in J,\alpha\in P}f(\delta_{\alpha_{j}})\mu(\alpha_{j})|\leq u_{n}\mu(\Omega\setminus N)+\sum_{\alpha\in P}f_{n}(\delta_{\alpha})\mu(\alpha).$
Therefore, the net $\{\sum_{j\in J,\alpha\in P}f(\delta_{\alpha_{j}})\mu(\alpha_{j})\}_{J\in\mathcal{F}(\mathbb{N})}$ is bounded for the given $P$, which shows, as $X$ is Dedekind complete, that the order limit of this net, which is given by:
$\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha),$
is a well defined quantity (i.e, the order limit exists). By $(ii)$ in the conventions above, this limit can also be taken in the sense of $(D)$-convergence. The restriction put on $P\in\mathcal{P}(\Omega\setminus N)$ is not so restrictive, as we will only need partitions for which the $S^{*}$-partition integrals of $f_{n}$ are well defined, thus fulfilling the argument above.
We repeat the first part above with different notation to avoid confusion: as $f_{n}$ converges to $f$ uniformly almost everywhere in order, there exists a non-increasing sequence $\{u_{k}\}_{k\in\mathbb{N}}$ such that $u_{k}\downarrow 0$ and:
$|f_{k}(x)-f(x)|\leq u_{k},\ \forall x\in\Omega\setminus N\ \text{and}\ k\in\mathbb{N},$ (10)
which implies that $\{f_{n}\}_{n\in\mathbb{N}}$ is Cauchy uniformly.
Therefore, by Proposition 4.4, there exists $L\in Z$ such that the following order limit is valid:
$\int^{S^{*}}_{\Omega}f_{n}d\mu=\int^{S^{*}}_{\Omega\setminus N}f_{n}d\mu\xrightarrow[]{o}L.$
In other words, there exists a non-decreasing sequence $\{w_{n}\}_{n\in\mathbb{N}}\in X$ such that $w_{n}\downarrow 0$ and:
$|\int^{S^{*}}_{\Omega}f_{n}d\mu-L|\leq w_{n},\ \forall n\in\mathbb{N}.$ (11)
As each $f_{n}$ is $S^{*}$-partition integrable in on $\Omega\setminus N$, for each $n\in\mathbb{N}$, there exists a $(D)$-sequence $\{a^{(n)}_{ij}\}_{ij}\in Z$ such that, $\forall\ \psi\in\mathbb{N}^{\mathbb{N}}$, there exists $P_{n}=P_{n}(\psi)\in\mathcal{P}(\Omega\setminus N)$ such that for each $P\geq P_{n}$ in $\mathcal{P}(\Omega\setminus N)$ the unconditionally convergent series of $f_{n}$ and $\mu$ exists in $P$ and:
$|\sum_{\alpha\in P}f_{n}(\delta_{\alpha})\mu(\alpha)-\int^{S^{*}}_{\Omega}f_{n}d\mu|\leq\bigvee_{i=1}^{\infty}a^{(n)}_{i\psi(i)}$ (12)
As $\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)$ exists, by the argument in the first part of the present proof, for $P\geq P_{n}$, using (10),(11) and (12), the following inequalities are valid:
$|\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)-L|\leq|\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)-\sum_{\alpha\in P}f_{n}(\delta_{\alpha})\mu(\alpha)|$
$|\sum_{\alpha\in P}f_{n}(\delta_{\alpha})\mu(\alpha)-\int^{S^{*}}_{\Omega}f_{n}d\mu|+|\int^{S^{*}}_{\Omega}f_{n}d\mu-L|$
$\leq|\sum_{\alpha\in P}(f(\delta_{\alpha})-f_{n}(\delta_{\alpha}))\mu(\alpha)|+\bigvee_{i=1}^{\infty}a_{i\phi(i)}+w_{n}.$
Now, for each $P$ as above, the net $\{\sum_{j\in J,\alpha\in P}|f_{n}(\delta_{\alpha_{j}})-f(\delta_{\alpha_{j}})|\mu(\alpha_{j})\}_{J\in\mathcal{F}(\mathbb{N})}$ is order and $(D)$- convergent to its supremum, as it is increasing and bounded above by $u_{n}\mu(\Omega\setminus N)$ for each $n$ and $X$ is Dedekind complete. Therefore,
$\sum_{j\in J,\alpha\in P}|f_{n}(\delta_{\alpha_{j}})-f(\delta_{\alpha_{j}})|\mu(\alpha_{j})\leq\sum_{\alpha\in P}|f_{n}(\delta_{\alpha})-f(\delta_{\alpha})|\mu(\alpha_{j}),$
for each $J\in\mathcal{F}(\mathbb{N})$. Hence, by Proposition 1.20 of [44]:
$|\sum_{\alpha\in P}f_{n}(\delta_{\alpha})-f(\delta_{\alpha})\mu(\alpha_{j})|\leq\sum_{\alpha\in P}|f_{n}(\delta_{\alpha})-f(\delta_{\alpha})|\mu(\alpha_{j}),$
Therefore, from the last main inequality:
$\leq\sum_{\alpha\in P}|f_{n}(\delta_{\alpha})-f(\delta_{\alpha})|\mu(\alpha)+\bigvee_{i=1}^{\infty}a^{(n)}_{i\psi(i)}+w_{n}$
$\leq\sum_{\alpha\in P}u_{n}\mu(\alpha)+\bigvee_{i=1}^{\infty}a^{(n)}_{i\psi(i)}+w_{n}=u_{n}\mu(\Omega\setminus N)+\bigvee_{i=1}^{\infty}a^{(n)}_{i\psi(i)}+w_{n}$
In resume, until now we have the following inequality:
$|\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)-L|\leq u_{n}\mu(S)+w_{n}+\bigvee_{i=1}^{\infty}a^{(n)}_{i\psi(i)},\ \forall\ P\geq P_{n}.$ (13)
Now, consider the following equation defining a triple sequence $\{c_{nij}\}_{nij}$ in $Z$:
$c_{nij}=\begin{cases}w_{j}+u_{j}\mu(\Omega\setminus N),&\text{if $n=1$},\\ a^{(n-1)}_{ij},&\text{if $n\geq 2,$}\end{cases}$
for $i,j\in\mathbb{N}$.
In this case, notice that, for each $n\in\mathbb{N}$ fixed, the (bounded) double sequence $\{c_{nij}\}_{ij}$ is a $(D)$-sequence of elements of $Z$ because, for each $i\in\mathbb{N}$, $\{c_{nij}\}_{j\in\mathbb{N}}$ is non-increasing and $c_{nij}\downarrow 0$, when $j\rightarrow\infty$.
Now, let $L^{{}^{\prime}}=u_{1}\mu(\Omega\setminus N)+w_{1}+\bigvee_{i,j=1}^{\infty}a^{(1)}_{ij}$. If $L^{{}^{\prime}}=0$, then $u_{1}=0$ and, by (10), $f=f_{k}$ $\forall k\in\mathbb{N}$ outside of the null set $N$, which finishes the proof. Assuming, then, that $L^{{}^{\prime}}\in Z^{+}\setminus\{0\}$ exists, by Fremlin Lemma, there exists a double sequence $\{b_{ij}\}_{ij}$ in $X$ such that $\{L^{{}^{\prime}}\wedge b_{ij}\}_{ij}$ is a $(D)$-sequence and, $\forall\psi\in\mathbb{N}^{\mathbb{N}}$,
$L^{{}^{\prime}}\wedge(\sum_{r=1}^{n}\bigvee_{i=1}^{\infty}c_{ri\psi(i+r)})\leq\bigvee_{j=1}^{\infty}(L^{{}^{\prime}}\wedge b_{j}\psi(j)),\ \forall n\in\mathbb{N}.$ (14)
We now show, that, the $(D)$-sequence $\{a_{ij}\}_{ij}$ given by:
$a_{ij}=(L^{{}^{\prime}}\wedge b_{ij}),$
is the $(D)$-sequence for the existence of the $S^{*}$-partition integral of $f$ on $\Omega\setminus N$.
In fact, given $\varphi\in\mathbb{N}^{\mathbb{N}}$, let $p=\min_{i\in\mathbb{N}}\varphi(i+1)$ and $\psi\in\mathbb{N}^{\mathbb{N}}$ the function defined by $\psi(j)=\varphi(j+p+1)$, $\forall j\in\mathbb{N}$. By (13), for $P\geq P_{p}$, $P_{p}=P_{p}(\varphi)$, all in $\mathcal{P}(\Omega\setminus N)$, we have that:
$|\sum_{\alpha\in P}f(\delta_{\alpha})-L|\leq u_{p}\mu(\Omega\setminus N)+w_{p}+\bigvee_{i=1}^{\infty}a^{(p)}_{i\varphi(i+p+1)}.$ (15)
Also, we have:
$\bigvee_{j=1}^{\infty}(u_{\varphi(j+1)}\mu(\Omega\setminus N)+w_{\varphi(j+1)})=u_{p}\mu(\Omega\setminus N)+w_{p}.$
Therefore, from (15), we may write that:
$|\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)-L|\leq\bigvee_{j=1}^{\infty}(u_{\varphi(j+1)}\mu(\Omega\setminus N)+w_{\varphi(j+1)})+\bigvee_{i=1}^{\infty}a^{(p)}_{i\varphi(i+p+1)},$
that is,
$|\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)-L|\leq\bigvee_{j=1}^{\infty}c_{ij\varphi(j+1)}+\bigvee_{i=1}^{\infty}c_{(p+1)i\varphi(i+p+1)}\leq\sum_{r=1}^{p+1}\bigvee_{i=1}^{\infty}c_{ri\phi(i+r)}.$ (16)
On the other hand, by (13), and for $n=1$, we have that:
$|\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)-L|\leq u_{1}\mu(\Omega\setminus N)+w_{1}+\bigvee_{i,j=1}^{\infty}a^{(1)}_{ij}=L^{{}^{\prime}},$ (17)
whenever $P\geq P_{1}$, with, as above, $P_{1}\in\mathcal{P}(\Omega\setminus N)$.
Now, taking $P_{0}=P_{1}\wedge_{\mathcal{P}(\Omega\setminus N)}P_{p}$, the refinement of $P_{1}$ and $P_{p}$, for $P\geq P_{0}$, we have that, by (14), (16) and (17),
$|\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)-L|\leq L^{{}^{\prime}}\wedge(\sum_{r=1}^{p+1}\bigvee_{i=1}^{\infty}c_{ri\varphi(i+r)})\leq\bigvee_{j=1}^{\infty}(L^{{}^{\prime}}\wedge b_{j\varphi(j)})=\bigvee_{j=1}^{\infty}a_{j\varphi(j)},$
which shows that, as indicated before, that $\{a_{ij}\}_{ij}$ is indeed the $(D)$-sequence for the existence of the $S^{*}$partition integral of $f$ on $\Omega\setminus N$, implying that $f$ is $S^{*}$-partition integrable on $\Omega\setminus N$ and by an appeal to Proposition 2.5:
$\int^{S^{*}}_{\Omega}f_{n}d\mu\rightarrow\int^{S^{*}}_{\Omega\setminus N}fd\mu,$
This finishes the proof.
We now prove a result to connect the convergence theorems related to the $S^{*}$-partition integral, outside of null sets, to the whole $\Omega$.
Theorem 4.3.
Suppose that $Z$ is Dedekind complete and let $f$ be a non-negative function $S^{*}$-partition integrable on $\Omega\setminus N$, where $N$ is a $\mu$-null set, $\mu:\mathcal{H}\rightarrow Y$ non-negative and additive. Then, $f$ is $S^{*}$-partition integrable on $\Omega$ and:
$\int^{S^{*}}_{\Omega}fd\mu=\int^{S^{*}}_{\Omega\setminus N}fd\mu$
Proof.
By hypothesis, following the notation in the theorem, for all choice functions $\delta$, there exists an $s\in Z$ and a $(D)$-sequence $\{a_{ij}\}$ such that, for every $\varphi\in\mathbb{N}^{\mathbb{N}}$, there exists $P_{0}\in\mathcal{P}(\Omega\setminus N)$ such that:
$|\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)-s|\leq\bigvee_{i=1}^{\infty}a_{i\varphi(i)},$
[…]
Consider now $P_{0}^{{}^{\prime}}=P_{0}\cup\{N\}$, which is an element of $\mathcal{P}(\Omega)$ by construction. Take now any $P^{{}^{\prime}}\geq P_{0}{{}^{\prime}}$ in $\mathcal{P}(\Omega)$. Define the following in relation to $P^{{}^{\prime}}$:
$A_{P^{{}^{\prime}}}=\{\alpha\in P^{{}^{\prime}}:\alpha\subseteq\beta,\text{for some}\ \beta\in P_{0}\},$
which satisfies, by definition:
$\bigcup_{\alpha\in A_{P^{{}^{\prime}}}}\alpha=\bigcup_{\beta\in P_{0}}(\bigcup_{\alpha\in P^{{}^{\prime}},\alpha\subseteq\beta}\alpha)=\bigcup_{\beta\in P_{0}}\beta=\Omega\setminus N,$
which gives that $A_{P^{{}^{\prime}}}\in\mathcal{P}(\Omega\setminus N)$ and $A_{P^{{}^{\prime}}}\geq P_{0}$.
As $N$ is a $\mu$-null set, we have that, for any $J\in\mathcal{F}(\mathbb{N})$,
$\sum_{j\in J,\alpha\in P^{{}^{\prime}}}f(\delta_{\alpha_{j}})\mu(\alpha_{j})=\sum_{j\in J,\alpha\in A_{P^{{}^{\prime}}}}f(\delta_{\alpha_{j}})\mu(\alpha_{j}),$
but $A_{P^{{}^{\prime}}}\in\mathcal{P}(\Omega\setminus N)$ and $A_{P^{{}^{\prime}}}\geq P_{0}$, therefore, by the integrability condition on $f$ enounced at the first paragraph of the proof, we get that the series
$\sum_{\alpha\in A_{P^{{}^{\prime}}}}f(\delta_{\alpha})\mu(\alpha),$
exists and as $f$ and $\mu$ are non-negative:
$\sum_{j\in J,\alpha\in P^{{}^{\prime}}}f(\delta_{\alpha_{j}})\mu(\alpha_{j})\leq\sum_{\alpha\in A_{P^{{}^{\prime}}}}f(\delta_{\alpha})\mu(\alpha),$
for every $J\in\mathcal{F}(\mathbb{N})$. As $Z$ is Dedekind complete,
$\sum_{\alpha\in P^{{}^{\prime}}}f(\delta_{\alpha})\mu(\alpha),$
exists for each $P^{{}^{\prime}}$ given above. By the same reasoning above, we have not only that the series index on $P^{{}^{\prime}}$ exists, but also:
$\sum_{\alpha\in P^{{}^{\prime}}}f(\delta_{\alpha})\mu(\alpha)=\sum_{\alpha\in A_{P^{{}^{\prime}}}}f(\delta_{\alpha})\mu(\alpha),$
and using the fact that $A_{P^{{}^{\prime}}}\geq P_{0}$, we get by the relation involving $s$,
$|\sum_{\alpha\in P^{{}^{\prime}}}f(\delta_{\alpha})\mu(\alpha)-s|\leq\bigvee_{i=1}^{\infty}a_{i\varphi(i)},$
which shows that $f$ is integrable on $\Omega$ and:
$s=\int^{S^{*}}_{\Omega\setminus N}fd\mu=\int^{S^{*}}_{\Omega}fd\mu,$
which finishes the proof.
Combining Theorems 4.3 and 4.2 we get the final result:
Theorem 4.4 (Almost Everywhere Uniform Converge Theorem for the $S^{*}$-partition Integral).
Let $X$ be a Dedekind complete Riesz space that is weakly $\sigma$-distributive and $\{f_{n}\}_{n\in\mathbb{N}}$ a sequence of non-negative $S^{*}$-partition integrable functions on $\Omega$. Suppose that $\mu:\mathcal{H}\rightarrow Y^{+}$ is $\sigma$-additive. If $f_{n}$ converges uniformly almost everywhere in order to $f$, a non-negative function, then $f$ is $S^{*}$-partition integrable on $\Omega$ and:
$\int^{S^{*}}_{\Omega}f_{n}d\mu\xrightarrow[]{o}\int^{S^{*}}_{\Omega}fd\mu,$
A almost identical proof (in fact, simpler one) can be made for the Sion integral, when considered with respect to partitions and not coverings, of all the results above for the $S^{*}$-partition integral related to null sets, which leads us to enounce:
Theorem 4.5 (Almost Everywhere Uniform Converge Theorem for the Sion Integral).
Let $X$ be a Dedekind complete Riesz space that is weakly $\sigma$-distributive, $\mu:\mathcal{H}\rightarrow Y^{+}$ a regular integrator, and $\{f_{n}\}_{n\in\mathbb{N}}$ a sequence of Sion (partition) integrable functions on $\Omega$. If $f_{n}$ converges uniformly almost everywhere in order to $f$ then $f$ is Sion (partition) integrable on $\Omega$ and:
$\int^{Si}_{\Omega}f_{n}d\mu\xrightarrow[]{o}\int^{Si}_{\Omega}fd\mu,$
We now have finished the convergence theorems involving the Net Riemann integrals and correlated concepts. An analogue of the General Convergence Theorem of [11] can be proved almost identically as in the cited reference. Analogues of the Henstock Lemma (see, for exemple, [67]) and a restricted type of monotone convergence theorem can be proved for the Net Riemann integral (or the $S^{*}$-partition integral) with further hypothesis, specially by using the connection with the derivation basis integral of [69] and [80], or the (particular case of our integration theory) integral of [54]. We leave these as open problems for a future work.
5 Pavlakos Lebesgue Type Integration and its Riemmanian Properties
We now expound some aspects of a very specific type of Lebesgue integral, by which we mean integrals defined by approximating the functions to be integrate by a net of simple/elementary functions, for functions and measures with values in Riesz spaces. For a brief and incomplete survey of such constructions, we mention the uniform net limit type integrals of [81], [30], [31], [29], the sequential integrals of [32], [82], [83], [84], [85], [86], [87], [88], [89], [90], [91], [92], [93], [94], [95], [96], [97], [98] [38], [99], [100], [57], [12], [39], [101], [102], [103], [104], [105], [42], [43], [41] the theses of [106], [107], [108], [109], containing abundant information on the structures that will be discussed here, as well as the related Daniell type integrals, not covered here in any detail, of [33], [52], [110], [111], [112], [113], [114], [115], [116], [117], [118], [119], [120], [121], [122], [123], [124], [125], [126], [127], [128], [129], [130], and [131].
Some of the constructions given above are valid, and in fact presented in the works of the original authors, for lattice ordered groups or partially ordered spaces more generally, but to maintain certain uniformity, and also apply the convergence theorems in the previous sections, we will put some restrictions on the $X,Y,Z$ and their convergence structures.
More specifically, we will restrict ourselves in this section to the following setting:
  1. 1.
    Riesz spaces $X,Y,Z$ such that $Z$ is Dedekind complete and weakly $\sigma$-distributive. In relation to the measure theoretic objects, $(\Omega,\mathcal{H})$ will always denote a measurable space, i.e, $\mathcal{H}$ is a $\sigma$-algebra of subsets of $\Omega$, and $\mu:\mathcal{H}\rightarrow Y^{+}$ will be $\sigma$-additive in the net convergence structure of $Y$ in the sense of Definition 3.13. We also suppose that $\mu$ is a regular integrator.
  2. 2.
    In relation to convergence, we shall equip $Z$ with the net convergence structure of $(D)$-convergence or order convergence (each case will be explicitly specified), and $X,Y$ with order convergence in the sense of Definition 2.3 (which, in this case, is equivalent to order convergence in the sense of Definition 2.2).
Further hypothesis of any sort will be pointed out when necessary.
With these observations, we pass to a brief presentation the integration theory of Pavlakos [30], which contains the integral of Kappos [29] and Berg [81].
To develop his integral, Pavlakos [30] supposes that $X$ is of countable type with respect to order convergence, i.e he supposes $X$ is fist countable with respect to the net convergence structure given by order convergence. Therefore, we will also suppose throughout this section, besides 1. and 2. above, that:
3. $X$ is first countable with respect to the net convergence structure of order convergence.
Now, we define a list of functions spaces specified by Pavlakos in different sections in [30]. To obtain clarity and brevity, we will define the function spaces in question already with the hypothesis put onto them by Pavlakos [30] for the purpose of integration, and with the same notation.
Definition 5.1 (Function Spaces).
For $X,Y,Z$ and $(\Omega,\mathcal{H})$ fixed as above, we define:
  1. (i)
    $F(\Omega,X)$, the space of all functions $f:\Omega\rightarrow X$ equipped with addition and multiplication by (real) scalars pointwise, and with the pointwise partial order inherited from $X$. In this case, it becomes a Riesz Space.
  2. (ii)
    $E(\Omega,X)$, the (Riesz subspace) of $F(\Omega,X)$ consisting of all $\mathcal{H}$-elementary functions, i.e: the space of all functions $f$ of the form
    $f(\omega)=\sum_{n=1}^{\infty}a_{i}\mathbbm{1}(\omega)_{A_{i}},$
    where $\{a_{n}\}_{n=1}^{\infty}$ is a sequence of elements of $X$ and $\{A_{n}\}_{n\in\mathbb{N}}$ is a sequence of disjoint sets from $\mathcal{H}$.
  3. (iii)
    $M(\Omega,X)$, the (Riesz subspace) of $F(\Omega,X)$ consisting of all functions $f$ such that there exists a sequence $\{f_{n}\}_{n\in\mathbb{N}}\in E(\Omega,X)$ such that
    $f_{n}\rightarrow f,$
    uniformly almost everywhere in order as defined in the previous section, or $(u)-\lim_{n\rightarrow\infty}f_{n}=f$ a.e.-$\mu$ on $\Omega$ in [30] notation and terminology. This is called the space of $(\mathcal{H},\mu)$-measurable functions.
We now define the integral in question by steps, as is usual for Lebesgue type integration procedures. For a non-negative function $f\in E(\Omega,X)$ of the form
$f=\sum_{n=1}^{\infty}a_{i}\mathbbm{1}_{A_{i}},$
with the same notation as in the previous definition, we say that it is Pavlakos integrable if the following (unconditional series!) limit:
$(o)-lim_{n\rightarrow\infty}\sum_{i=1}^{n}a_{i}\mu(A_{i}),$
exists in $Z$. Then, in this case we use the notation:
$\int_{\Omega}^{P}fd\mu=(o)-lim_{n\rightarrow\infty}\sum_{i=1}^{n}a_{i}\mu(A_{i})=\sum_{n=1}^{\infty}a_{i}\mu(A_{i}).$
By Lemmas 1.2 and 3.1 of [30], this quantity is well defined in the sense of not depending on the rearrangement of the series and the representation of $f$ as an elementary function (which is, in general, non-unique).
For a general $\in E(\Omega,X), ff$ is Pavlakos integrable if $f^{+}$ and $f^{-}$ are integrable. In this case, we define:
$\int_{\Omega}^{P}fd\mu=\int_{\Omega}^{P}f^{+}d\mu-\int_{\Omega}^{P}f^{-}d\mu,$
with the integral in a set $A\in\mathcal{H}$ being defined as:
$\int_{A}^{P}fd\mu=\int_{\Omega}^{P}f\mathbbm{1}_{A}d\mu.$
By Theorem 3.5 and Corollary 3.7 of [30], the defined integral on $E(\Omega,T)$ is a linear, positive and isotone operator and it is absolute, in the sense that $f$ is Pavlakos integrable if and only if $|f|$ is Pavlakos integrable.
We now enounce Theorems 3.11 and Corollary 3.12 of [30], respectively, for the definition of the integral on $M(\Omega,X)$:
Theorem 5.1.
Let $\{f_{n}\}_{n\in\mathbb{N}}$ be an increasing sequence of $E(\Omega,T)$ Pavlakos integrable functions.
Let $f\in M(\Omega,X)$ such that:
$(u)-\lim_{n\rightarrow\infty}f_{n}=f\ \mu-\text{a.e. on}\ \Omega.$
Then, there exists the order limit $(o)-\lim_{n\rightarrow\infty}\int_{\Omega}f_{n}d\mu$.
Lemma 5.1.
Let $\{f_{n}\}_{n\in\mathbb{N}}$ and $\{b_{n}\}_{n\in\mathbb{N}}$ be increasing sequences of Pavlakos integrable functions on $E(\Omega,X)$ such that:
$(u)-\lim_{n\rightarrow\infty}f_{n}=(u)-\lim_{n\rightarrow\infty}g_{n},\mu-\text{a.e. on}\ \Omega.$
Then,
$(o)-\lim_{n\rightarrow\infty}\int^{P}_{\Omega}f_{n}d\mu=(o)-\lim_{n\rightarrow\infty}\int^{P}_{\Omega}g_{n}d\mu.$
We then have the main integral of [30]:
Definition 5.2 (Pavlakos Integral).
A function $f\in M(\Omega,X)$ is said to be $(\mathcal{H},\mu)$-Pavlakos integrable on $\Omega$ if there exists an increasing sequence $\{f_{n}\}_{n\in\mathbb{N}}$ in $E(\Omega,X)$ of Pavlakos integrable functions such that $(u)-\lim_{n\rightarrow\infty}f_{n}=f$ $\mu$-a.e. on $\Omega$.
In this case, we the integral of $f$ on $\Omega$ is the element of $Z$ defined by the equality
$\int^{P}_{\Omega}fd\mu=(o)-\lim_{n\rightarrow\infty}\int^{P}_{\Omega}f_{n}d\mu$
It is a consequence of Theorem 5.1 and Lemma 5.1 that this integral is well defined. The definition of this integral on a set $A\in\mathcal{H}$ and its properties are as same as before. For more basic properties, we refer the reader to the original article of Pavlakos [30].
The last special property of this integral that we will need is the following Theorem 3.8 of [30]:
Theorem 5.2 ($\sigma$-Additivity of the Pavlakos Integral).
Let $f$ be a non-negative Pavlakos integrable function, then the set function $\nu:\mathcal{H}\rightarrow Z^{+}$ defined by:
$\nu(A)=\int_{A}^{P}fd\mu,$
is $\sigma$-additive in $\mathcal{H}$.
We now begin to present the connections of this integral with the previous ones of Riemann type. To obtain such a comparison, consider the following result:
Proposition 5.1.
Let $f:\Omega\rightarrow X^{+}$ be a non-negative elementary function of the form:
$f=\sum_{n=1}^{\infty}a_{i}\mathbbm{1}_{A_{i}},$
which is Pavlakos integrable on $\Omega$ with:
$\int_{\Omega}^{P}fd\mu=(o)-\lim_{n\rightarrow\infty}\sum_{i=1}^{\infty}a_{i}\mu(A_{i})=\sum_{n=1}^{\infty}a_{n}\mu(A_{n}).$
Then, for every $P\in\mathcal{P}(\bigcup_{n=1}^{\infty}A_{n})$ of the form $P=\{B_{m}\}_{m=1}^{\infty}$ such that $P\geq P_{0}=\bigcup_{n=1}^{\infty}\{A_{n}\}$, then:
$\sum_{m=1}^{\infty}b_{i}\mu(B_{i})=\sum_{n=1}^{\infty}a_{i}\mu(A_{i}),$
where $b_{i}$ is the value of $f$ at $B_{i}$.
Proof.
Let $P,P_{0}$ be as in the hypothesis of the proposition. Then, by definition:
$\bigcup_{m=1}^{\infty}B_{m}=\bigcup_{n=1}^{\infty}A_{n}.$
Therefore, as $f$ is non-negative and Pavlakos Integrable on $\Omega$, by Theorem 5.2, we have that:
$\int_{\bigcup_{m=1}^{\infty}B_{m}}^{P}fd\mu=\sum_{m=1}^{\infty}\int_{B_{m}}^{P}fd\mu=\sum_{n=1}^{\infty}a_{n}\mu(A_{n}).$
But, as $P$ refines $P_{0}$, for each $B_{i}\in P$ there exists (a unique) $A_{j}\in P_{0}$ such that $B_{i}\subseteq A_{j}$. Therefore, $f$ restricted to $B_{i}$, $f|_{B_{i}}$, is such that:
$f|_{B_{i}}:=b_{i}=a_{j},$
which implies that $f$ is constant on each element of $P$ and thus from the definition of the Pavlakos integral we get that:
$\sum_{m=1}^{\infty}\int_{B_{m}}^{P}fd\mu=\sum_{m=1}^{\infty}b_{m}\mu(B_{m}),$
from which we get:
$\sum_{m=1}^{\infty}b_{m}\mu(B_{m})=\sum_{n=1}^{\infty}a_{n}\mu(A_{n}),$
which finishes the proof.
From this proposition, we get:
Corollary 5.2.1.
Let $f:\Omega\rightarrow X^{+}$ be a non-negative elementary function of the form:
$f=\sum_{n=1}^{\infty}a_{i}\mathbbm{1}_{A_{i}},$
which is Pavlakos integrable on $\Omega$ with:
$\int_{\Omega}^{P}fd\mu=(o)-\lim_{n\rightarrow\infty}\sum_{i=1}^{\infty}a_{i}\mu(A_{i})=\sum_{n=1}^{\infty}a_{n}\mu(A_{n}).$
Then, $f$ is $S^{*}$-partition on $\Omega$ and:
$\int_{\Omega}^{S^{*}}fd\mu=\sum_{n=1}^{\infty}a_{n}\mu(A_{n})$
Proof.
Without loss of generality, suppose that $\Omega=\bigcup_{n=1}^{\infty}A_{n}$, and take $P_{0}=\bigcup_{n=1}^{\infty}\{A_{n}\}$. Let $P\in\mathcal{P}(\Omega)$ such that $P\geq P_{0}$. By construction (and the same argument as in the proof of the previous proposition), for each choice function $\delta$, $f(\delta_{\alpha})$, $\alpha\in P$, is a constant, and:
$\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha):=\sum_{m=1}^{\infty}b_{m,P}\mu(B_{m,P}),$
where $P=\bigcup_{m=1}^{\infty}\{B_{m,P}\}$, the subindex $m,P$ indicates that the sets corresponds to the partition $P$, and for which the existence of the (unconditional) series in the right-hand side follows from Proposition 5.1. In fact, from the same proposition, we get that:
$\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)=\sum_{m=1}^{\infty}b_{m,P}\mu(B_{m,P})=\sum_{n=1}^{\infty}a_{n}\mu(A_{n}).$
[…]
$|\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)-\sum_{n=1}^{\infty}a_{n}\mu(A_{n})|\leq\bigvee_{i=1}^{\infty}a_{i\varphi(i)},$
which shows that $f$ is $S^{*}$-partition integral with the same value as the Pavlakos integral of $f$. ∎
We now get to the main comparison result:
Theorem 5.3 (Pavlakos Integrability implies $S^{*}$-partition Integrability - Non-negative Case).
Let $f:\Omega\rightarrow X$ be a non-negative Pavlakos integrable function on $\Omega$. Then, $f$ is $S^{*}$-partition integrable on $\Omega$ and the two integrals agree.
Proof.
By the definition of the Pavlakos integrability of $f$, let $\{f_{n}\}_{n\in\mathbb{N}}$ be a sequence of non-negative increasing functions, in $E(\Omega,X)$, of Pavlakos integrable functions such that $(u)-\lim_{n\rightarrow\infty}f_{n}=f$ $\mu$-a.e. on $\Omega$.
By Corollary 5.2.1, we get that each $f_{n}$ is $S^{*}$-partition integrable and:
$\int_{\Omega}^{P}fd\mu=\int_{\Omega}^{s^{*}}fd\mu.$
On the other hand, by the almost everywhere uniform convergence Theorem for the $S^{*}$-partition integral, Theorem 4.4, we have that $f$ is $S^{*}$-partition integrable on $\Omega$:
$(o)-\lim_{n\rightarrow\infty}\int_{\Omega}^{S^{*}}f_{n}d\mu=\int_{\Omega}^{S^{*}}fd\mu.$
As we have that, by definition:
$(o)-\lim_{n\rightarrow\infty}\int_{\Omega}^{P}f_{n}d\mu=\int_{\Omega}^{P}fd\mu,$
it follows that:
$\int_{\Omega}^{P}fd\mu=\int_{\Omega}^{S^{*}}fd\mu,$
which finishes the proof. ∎
For the general case, we simply write the decomposition of a Pavlakos integrable $f$ on its negative and positive parts, both non-negative and Pavlakos integral, use Theorem 5.3, and then we get:
Corollary 5.3.1 (Pavlakos Integrability implies $S^{*}$-partition Integrability - General Case).
Let $f:\Omega\rightarrow X$ be a Pavlakos integrable function on $\Omega$. Then, $f$ is $S^{*}$-partition integrable on $\Omega$ and the two integrals agree.
Related arguments can be made for the Sion integral in the place of the $S^{*}$-partition integral, but they are similar, so we will not state them.
One interesting aspect of the discussion from now on is, then, to connect net-approximation (Lebesgue) type integrals to the Pavlakos’ one in an attempt to relate them to the Riemmanian approach. As real measurable non-negative functions $f:\Omega\rightarrow\mathbb{R}^{+}$ are strongly measurable in the sense of Böchner (see [132]), and thus can be approximated uniformly by a sequence of increasing elementary functions (see again [132]), we may use the monotone convergence theorem of Maitland-Wright [32] and the device of decomposing a function into negative and positive parts to prove:
Proposition 5.2.
Let $f:\Omega\rightarrow\mathbb{R}$ be a Maitland-Wright integrable function on $\Omega$, then it is Pavlakos integrable on $\Omega$ and the two integrals agree.
For more details on the Maitland-Wright integral, see [32] and [82], and [30]. A similar result can be obtained for the integral of Kusraev and Tasoev [102].
This result also helps to connect the Riemmanian approaches to integration more closely with the sequential/net-based Lebesgue integrals in the most well known approaches to integration in partially ordered spaces, at least for real functions. We now pass to a brief discussion of a general case for Lebesgue and Saks-type integration, and comment more about the still-to-be established results classifying all these integration procedures.
6 General Lebesgue and Saks Type Integration on Boccuto-Candeloro Net Convergence Structures
6.1 Abstract Lebesgue Integration
Taking the cues in the last section, we see that, in general, all the Lebesgue type procedures for integration cited previously follows the same methodology:
  1. (i)
    Choose a space of functions $L$ that contains the simple (measurable) functions and define the integral for this former class in some sense. For the simple functions we define the integral, generally, as a linear combination of the constants and measure of the sets in the definition.
  2. (ii)
    Choose now a class of functions that can be approximated, in the sequential or net sense, by the functions in $L$. Prove that the limit of the integrals of such functions is independent of the chosen approximation.
  3. (iii)
    Define, then, using the last point, the integral of a function $f$ as the limit of the integrals of the approximating functions on $L$.
The convergence properties given by a topology, order, or other type of (net/filter) convergence structure appears in a very essential way in point (ii) - most of the time to produce uniqueness of the integral for each approximating sequence. In this sense, this strategy for obtaining an integration theory is dependent on the details of the convergence involved, being difficult to generalize to arbitrary convergence structures, something that can be done, as shown above, for the Riemmanian case.
Nevertheless, a kind of general Lebesgue integral can be propose in this section if we let the uniqueness portion of the procedure be a part of the definition of this integral - something that can be checked in every case of interest (e.g, the references above).
This general integral is based on an extension of the sequential definition given in Section 3.2 of [12], and also in the ideas of Boccuto and Sambucini in [133] and [39].
In this subsection, we consider the following setting:
  1. 1.
    $(X,\eta_{1}),(Y,\eta_{2}),(Z,\eta_{3})$ three spaces equipped with Boccuto-Candeloro net convergence structures, and connected by a product as in Definition 3.6,
  2. 2.
    $(\Omega,\mathcal{H})$ an abstract paved space,
  3. 3.
    $\mu:\mathcal{H}\rightarrow Y^{+}$ a finitely additive set function.
Now, to begin the procedures to define the integral, consider a space of functions $L$, consisting of all $f:\Omega\rightarrow X$ satisfying some fixed property (e.g, being simple, bounded, etc) such that it contains all simple functions defined on $(\Omega,\mathcal{H})$, and with the property that $f\mathbbm{1}_{A}\in L$, for each $A\in\mathcal{H}$.
Suppose now that there exists a functional denoted by:
$\int_{\Omega}^{L}(\cdot)\ d\mu:L\rightarrow Z,$
such that:
  1. 1.
    $\int^{L}_{\Omega}(\cdot)\ d\mu$ is linear,
  2. 2.
    $\int^{L}_{\Omega}(\cdot)\ d\mu$ is such that, for each simple function $g=\sum_{i=1}^{n}a_{i}\mathbbm{1}_{A}$,
    $\int^{L}_{\Omega}gd\mu=\sum_{i=1}^{n}a_{i}\mu(A_{i}),$
Now put $\int^{L}_{A}fd\mu=\int^{L}_{\Omega}f\mathbbm{1}_{A}d\mu$.
Definition 6.1 (Abstract Lebesgue Integral).
A non-negative function $f:\Omega\rightarrow X$ is said to be Abstractly Lebesgue integrable, or just Lebesgue integrable in case no confusion arises, if there is a net $\{f_{i}\}_{i\in I}$ of functions in $L$ such that $f_{i}\rightarrow f$ in some convergence defined on the function space $L$ with respect to $(X,\eta_{1})$ [^7] For example: the uniform or almost uniform convergences defined in the previous sections., which we call the approximation procedure, and such that the limit:
$\lim_{(I,\eta_{3})}\int^{L}_{A}f_{i}d\mu,$
exists in $Z$ for every $A\in\mathcal{H}$, and is independent of the approximating net $\{f_{i}\}_{i\in I}$. In this case, we denote by the set function produced for $f$, and for each $A$ in the limit above, as $I_{f}(\cdot)$, and call:
$I_{f}(A)=\lim_{(I,\eta_{3})}\int^{L}_{A}f_{i}d\mu,$
the Abstract Lebesgue integral of $f$ on $A$.
In this case, an arbitrary function $f:\Omega\rightarrow X$ is called Abstractly Lebesgue integrable if $f^{+}$ and $f^{-}$ are and in this case the integral of $f$ is defined to be:
$I_{f}(A)=I_{f^{+}}(A)-I_{f^{-}}(A),$
where $A\in\mathcal{H}$.
The first thing to notice is that this integral is readily seen to be a linear operator on the space of Lebesgue integrable functions. Other properties of this integral seem to require special hypothesis, such as the uniform integrability property of the net approximation of functions as in [12] or some other restrictive property of the approximation procedure.
We also comment that not only the sequential (general) Lebesgue integral of [12] is contained in our definition, but also most examples of the Lebesgue type integrals cited in the beginning of Section 5. In particular, taking $L$ to be the space of elementary functions, and approximation procedure of almost uniform convergence, we get the Pavlakos Integral. Also, by analogous constructions, the recently defined integrals of [41], [42], defined on the basis of a sequential (monotone) approximation by simple functions, are also contained in our general definition. Lastly, by the results in Section 5, we get some some connections between this general Lebesgue Integral and the Riemmanian ones.
We stop the development of this integral at this point, and refer the reader to [12], [133] and [39] for more properties in some specific cases, and also point that some other questions will be answered in a future work.
6.2 A Saks Type Integral
Now, we pass to a brief discussion of the last type of integral to be defined in this work, a kind of Saks type integral as defined originally by Saks [134] and Gould [135], but also more recently by Fleischer [15] and Shcherbin [19]. Basically, the idea is to integrate functions by first integrating in an upwards directed class of sets, and then take the limit of integration in sets in this class as it "expands" (as it is (upwards) directed by inclusion, therefore forming a net). The idea is the same as in classical Lebesgue integration, where we may define the integral in an unbounded set as the limit of integrals on approximating (upwards) bounded sets.
For this section, we mainly follow the ideas of Fleischer [15] (which are, interestingly, presented in a convergence structure type language), and the same structures as in the Definition 3.9 of the Fleischer Net Riemann integral.
In this case, we have, by the same notation as in Section 3.1 of the (Fleischer) Net Riemann Integral:
Definition 6.2 (Saks Integral).
Let $f:\Omega\rightarrow X$ and $\mu:\mathcal{H}\rightarrow Y$ be arbitrary function and set function respectively. Also, suppose that, for each $A\in\mathcal{H}$, it is given an non-empty upwards directed by inclusion class of sets $\mathcal{C}_{f,A}$, which also depends on $f$, such that:
$\mathcal{C}_{A}\subseteq\mathcal{H},$
and $f$ is Fleischer Net Riemann integrable on each $B\in\mathcal{C}_{f,A}$. Then, we say that $f$ is Saks integrable on $A$ if the net of Fleischer Net Riemann integrals $\{\int_{B}fd\mu\}_{B\in\mathcal{C}_{f,A}}$ has a limit in $Z$ and, in this case, we define:
$\int_{A}^{Sa}fd\mu=\lim_{(\mathcal{C}_{f,A},\eta_{3})}\int_{B}fd\mu,$
and call this the Saks integral of $f$ on $A$.
As we suppose that the space $Z$ is Hausdorff in its Boccuto-Candeloro net convergence structure, the integral just presented is well defined (taking into account that $\mathcal{G}_{f,A}$ is directed). Taking also the Fleishcer Net Riemann integral gives more flexibility in defining the integral for each $A\in\mathcal{H}$, instead only for the whole space (as in [15]).
As in the previous subsection, we are not going to study this integral in detail, but just point some interesting aspects of it in connection to the pervious ones.
In fact, if $f$ is already Fleischer Net Riemann integrable on $A\in\mathcal{H}$, taking $\mathcal{G}_{f,A}=\{A\}$ produces the following:
Proposition 6.1.
If $f:\Omega\rightarrow X$ is Fleischer Net Riemann integrable on $A\in\mathcal{H}$, it is Saks integrable on $A$ and the two integrals agree.
Proof.
By the choice indicated above, i.e $\mathcal{G}_{f,A}=\{A\}$ (which is trivially upwards directed by inclusion), it follows that the net $\{\int_{B}fd\mu\}_{B\in\mathcal{G}_{f,A}}$ is constant and, by $(v)$ of Definition 3.3, its limit exists and equals the Fleischer Net Riemann integral $\int_{A}fd\mu$, which finishes the proof. ∎
Also, by ways of comparison, we point out that [15] defines a similar integral (which we call the Fleischer integral) [^8] Not to be confused with the Fleischer Net Riemann integral. based on the Kolmogorov S-integral with $\mathcal{H}$ being a ring and $\mathcal{G}_{f,A}$ being an ideal of sets of this ring and independent of $A$ (i.e, he takes it dependent only on $f$, and the only domain of integration being $\Omega$). In this case, we have:
Proposition 6.2.
Suppose that $\mathcal{H}$ is a ring of subsets of $\Omega$. For each $f:\Omega\rightarrow X$ and $A\in\mathcal{H}$, take $\mathcal{G}_{f,A}$ to be an ideal of sets. Then, if for each $B\in\mathcal{G}_{f,A}$, $f$ is Kolmogorov Integrable (See 3.2) on $B$, the Saks integral exists and coincides with the Fleischer integral.
We also note that, by the work of Boccuto and Riecan [9], we may obtain that the Saks integral contains also the Choquet integral of Riesz space-valued functions ([8]), and the Sipos integral for the case of non-negative functions. In fact, let $\mathcal{H}$ be a $\sigma$-algebra and $C:\mathcal{H}\rightarrow Y$ be a capacity (for this notion, see [8]). Fix $f:\Omega\rightarrow X=\overline{\mathbb{R}}^{+}$ a function such that:
$\Sigma^{f}_{t}=\{\omega\in\Omega:f(\omega)\geq t\}\in\mathcal{H},\forall\ t\in\mathbb{R},$
and define $u_{f}(t)=u(t)=C(\Sigma^{f}_{t})$. Note that $u:\mathbb{R}\rightarrow Y$ is a monotone function. In this context, we shall need the next result (see Section 3 of [8]):
Proposition 6.3.
Let $g:[a,b]\rightarrow Y$ be a monotone function with values in $Y$ that is supposed to be Dedekind complete and $a,b]\subseteq\mathbb{R}$ a finite interval. Then, $u$ is Kolmogorov $S$- integrable with respect to the Lebesgue measure.
Now, for each $f$ take, in the notation above, $\Omega=\overline{\mathbb{R}}^{+}$ with $\mathcal{H}$ the $\sigma$-algebra of Lebesgue measurable sets on $[a,b]$, and
$\mathcal{C}_{f,\overline{\mathbb{R}}^{+}}=\mathcal{C}_{\overline{\mathbb{R}}^{+}}=\{[0,a]\subseteq\mathbb{R}^{+}:a\in\mathbb{R}\}.$
Then, this collection is upwards directed by inclusion.
Considering the Net Riemann integral to be the Kolmogorov S-integral with respect to the Lebesgue measure as in the last proposition, we get that the net of integrals $\{\int_{0}^{a}u(t)dt\}_{[0,a]\in\mathcal{C}_{\overline{\mathbb{R}}^{+}}}$ is well defined an increasing. In case the limit exists, i.e the Saks Integral of $u$ exists on the positive extended real line, it is given by:
$\int_{\overline{\mathbb{R}}^{+}}^{Sa}u(t)dt=\sup_{[0,a]\in\mathcal{C}_{\overline{\mathbb{R}}^{+}}}\int_{0}^{a}u(t)dt=\sup_{a>0}\int_{0}^{a}u(t)dt,$
where the two last integrals are in the sense of the Kolmogorov S-integral, and which produced exactly the Choquet integral of $f$ (fixed above) in Definition 3.3 of [8] with respect to $C$. Therefore, we get, by using also Theorem 3.6 of [8]:
Proposition 6.4 (Saks includes Choquet and (Non-Negative) Sipos).
Let $f:\Omega\rightarrow X=\overline{\mathbb{R}}^{+}$ be a function with the properties above. Then, $f$ is Choquet integrable with respect to $C$ if and only if it is Sipos integrable with respect to $C$ and both integrals are equal to:
$\int_{\overline{\mathbb{R}}^{+}}^{Sa}u(t)dt,$
the Saks integral of $u$ over the extended real line with respect to the Lebesgue measure, formed by nets of Kolmogorov S-integrable functions.
This shows that this idea of Riemann integration derived from Saks’ original ideas is also fruitful for the case of non-additive set functions.
Some convergence theorems (without proofs) for this integral in a more restricted setting can be found in the article [20].
We finish now the presentation of this integral, and indeed of all others, and pass to the final comments.
7 Classification Schemes and (Various) Open Problems
The various integrals presented above all possesses strengths and weaknesses, specially with respect to which classes of functions and measures are admitted as possible ingredients of the procedure. As Kolmogorov [13] argued, a general definition of an integration procedure is unlikely to contain all existing ones, for various reasons. Nonetheless, the development above provides a reason to expect that each integration procedure we can find falls within one of the four categories now presented:
  1. 1.
    Saks type-Integration,
  2. 2.
    Abstract Lebesgue Integration,
  3. 3.
    Monotone Type, modeled on Choquet and Sipos, Integration,
  4. 4.
    Weak Type Integration, after Pettis [136] famous model.
Given the variety and diversity of integration procedures exhibited in the present work, such a classification may be premature, but can generate a way of organizing all the massive existing literature about integration theories. In this respect, an open problem, which is more of a research program carried forward already by such works as [4], [67], [60] and others, is:
  • 1.
    Is there a general relationship between the general procedures above ? More generally, can a "taxonomy" of integration procedures be produced ?
An affirmative response to both problems seems unlikely, specially as some integration procedures are not even comparable, but even this is a useful response to the general question: is there a "most general" procedure in each class ? And, if so, what are its limits ?
A related issue is the one of convergence: general (net or filter) convergence structures provides integration theory wide generality, but further restrictions need to be made to obtain stronger convergence and other type of theorems for this integrals.
This provides the second (in our view, interesting) problem:
  • 1.
    Is it possible to obtain (uniform) convergence theorems for the integrals above without requiring some specific convergence, only working with general properties of the filters/net involved, maybe with some minor additional condition to guarantee good continuity behaviors of the convergence ?
This is basically a question of the continuity of a (linear and isotone, sometimes) functional on spaces with general convergences and with specific properties generated by integration. That seems to require additional study of general convergence structures and many-valued topologies as is [137], [138], [139], [140] and references and further works therein. This seems to be related to the connection between integration in general convergence structures and Daniell integration in this context - continuity properties here are the key to obtain possible connections.
With respect to measures and functions, it seems directly possible, and in fact straightforward, to modify all the Riemmanian constructions above, not so much the Lebesgue ones, to interval functions in the sense of Henstock [67]. This provides, in particular, a non-linear version of the integration procedures above. The domain of the measures produces another challenge:
  • 1.
    Considering more general domains, such as orthoalgebras, difference posets, (D)-lattices and other structures presented in [141], which results above are preserved and, more interestingly, which convergence properties of measures and integrals are possible to prove ?
This also relates to the general integration theory of Kusraev and Malyugin [101], which may serve as a guideline, specially with respect to extension theorems of measures.
Nothing has been said with respect to weak type integrals of Pettis [136] type. It is not difficult to prove that the Fleischer Net Riemann and Saks Type Integrals (but not the $S^{*}$-integral!), are of the weak type as in Sipos [142] weak type order integral, which is defined in terms of an "order dual" of a Riesz Space. Related to duality in general convergence (partially ordered groups and vector spaces) as in Chapter 4 of [138], it arises the following question:
  • 1.
    Can a general weak-type Pettis integration be developed in partially ordered net convergence structures in such a way that this theory is compatible with Riemann and Lebesgue type-integration procedures in these general convergences ?
A comparison with and study of [143] and [144] should be made in this case.
Lastly, we pose the following problem:
  • 1.
    The spaces of integrable functions $L^{1}$, with respect to each integration procedures, can be characterized as partially ordered vector spaces. Which convergence and lattice structures turns them in a Riesz space with "good" (order separability, for example) properties ?
The references [145] and [41] can be a first step in this direction for general convergences.
More can be said about a host of other problems (product measures and Fubini theorems, for example) relating to general integration theory and its applications other areas, but we finish the article here, noting that much more can be said about this subject: both in its own right and regarding its relation to other areas of modern mathematics.
Acknowledgement
This study was funded in part by the Coordination for the Improvement of Higher Education Personnel - CAPES - Funding Code 001. The author would like to thank Professors Francesco Russo (ENSTA - Paris), Christophe Gallesco (UNICAMP), David Fremlin (Exeter) and Lech Drewnowski (Poznan) for interesting conversations and remarks regarding integration theory.